### v2

# Experimental validation of a linear momentum and bluff-body model for high-blockage cross-flow turbine arrays

Abstract
In a confined flow, the performance of a turbine and its near-wake fluid dynamics depend on the blockage ratio, defined as the ratio of the turbine projected area to the channel cross-sectional area. While blockage is understood to increase the power coefficient for turbine “fences” spanning a channel, most investigations at the upper range of practically-achievable blockage ratios have been theoretical or numerical in nature. Furthermore, while linear momentum actuator disk theory is frequently used to model turbines in confined flows, as confinement increases, the ability of this idealized model to describe performance and flow fields has not been established. In this work, the performance and near-wake flow field of a pair of cross-flow turbines are experimentally evaluated at blockage ratios from $30\%$ to $55\%$. The fluid velocity measured in the bypass region is found to be well-predicted by the open-channel linear momentum model developed by Houlsby et al. [1], while the wake velocity is not. Additionally, self-similar power and thrust coefficients are identified across this range of blockage ratios when array performance is scaled by the modeled bypass velocity following Whelan et al.’s [2] adaptation of the bluff-body theory of Maskell [3]. This result demonstrates that, despite multiple non-idealities, relatively simple models can quantitatively describe highly confined turbines. From this, an analytical method for predicting array performance as a function of blockage is presented. Overall, this work illustrates turbine performance at relatively high confinement and demonstrates the suitability of analytical models for predicting and interpreting their hydrodynamics.
Suggested keywords
preprint: es2025mar20_790
I Introduction
Flowing water in river and tidal channels is a promising source of renewable power. The efficiency of individual turbines or arrays deployed in these channels is influenced by how much of the channel cross-section they occupy. For a row of turbines, this is represented by the blockage ratio, defined as the ratio between the turbine projected area and the channel cross-sectional area:
$\beta=\frac{A_{\mathrm{turbines}}}{A_{\mathrm{channel}}}.$ (1)
As the blockage ratio increases, the turbines present greater flow resistance, but confinement from the channel boundaries leads to increased flow through the turbines relative to unconfined conditions. Consequently, as confinement increases, so does turbine efficiency [4, 2, 5, 6]. While blockage augments the performance of all turbine designs, cross-flow turbines—which have an axis of rotation perpendicular to the direction of inflow—are particularly well-suited to exploit these physics since their rectangular projected area matches the profile of natural and constructed channels.
Garrett and Cummins [4] were the first to establish the theoretical basis for turbine performance in confined flows by applying one-dimensional linear momentum actuator disk theory (LMADT). For relatively low Froude numbers, they showed that the maximum efficiency in confined flow was augmented by a factor of $(1-\beta)^{-2}$. Subsequent studies by Houlsby et al. [1] and Whelan et al. [2] relaxed the assumption of low Froude number by allowing the free surface to deform in response to momentum extraction. In parallel, Nishino and Willden [5, 7] expanded the conceptual framework with a two-scale LMADT model for turbine arrays that considered confinement effects from both the channel boundaries (“global” blockage) and the presence of neighboring turbines within the array (“local” blockage). Vogel et al. [6] later combined these approaches into a single model that incorporated both free surface deformation and inter-turbine proximity effects, and Dehtyriov et al. [8] extended the two-scale model of Nishino and Willden [5] to an arbitrary number of scales (i.e., arrays of turbine sub-arrays) using a fractal-like approach.
Blockage effects have been the focus of multiple experimental and numerical studies (Table 1). Across these, turbine dynamics are generally consistent and aligned with theory: for all types of turbines, as blockage increases, so do the rotor thrust loading, maximum efficiency, and optimal operating speed (i.e., tip-speed ratio or reduced frequency). Studies that examine the flow field near confined turbines [9, 10, 11, 12, 13, 14, 15] find that an increase in blockage accelerates the flow through and around the turbine as predicted by theory, as well as narrows the turbine wake. Gauthier et al. [16] and Kinsey and Dumas [17] examined how blockage effects change for various turbine types in asymmetric confinement (i.e., different combinations of lateral blockage and vertical blockage for a given $\beta$). The authors parameterized confinement asymmetry as the ratio between lateral blockage and vertical blockage (or its reciprocal, if a larger value), and found turbine performance to be invariant for values of confinement asymmetry less than three. The interplay between rotor geometry and blockage effects has also been explored for cross-flow turbines by Goude and Ågren [18] and for axial-flow turbines by Schluntz and Willden [13] and Abutunis and Menta [19]. These studies found that higher-solidity rotors benefit more from blockage – a consequence of how their higher thrust coefficients couple with channel-level energetics.
Several groups have investigated analytical blockage corrections—which adjust observed turbine performance in confined flow to predict the performance of the same turbine in unconfined flow—and evaluated those corrections through experiments [2, 9, 20, 21, 12, 22, 23, 24] or simulations [16, 17, 25, 24, 26]. Such methods generally proceed from Glauert [27] (who originally proposed a blockage correction for airplane propellers) by assuming that there exists some freestream velocity in unconfined flow that yields the same thrust and velocity through the rotor as in the confined condition. This unconfined freestream velocity is found through an analytical model, often LMADT [e.g., 28, 2, 23, 29]. However, blockage corrections based on LMDAT can yield non-physical results for “highly-loaded” turbines since sufficiently high thrust coefficients necessitate reversed flow in the turbine wake [30, 31]. This limitation has prompted investigation into blockage corrections that are applicable to a broader range of turbine solidities and operating conditions. For example, Whelan et al. [2] proposed a LMADT-based correction inspired by the bluff-body theory of Maskell [3], which, unlike Glauert’s theory, assumes that the thrust on a turbine depends on the velocity of the accelerated flow that bypasses it (i.e., the rotor thrust and bypass velocity are unchanged between corresponding confined and unconfined conditions). Whelan et al. found that such a correction more accurately predicted unconfined thrust coefficients for highly-loaded turbines. Recently, Steiros et al. [24] developed an alternative blockage correction for high-solidity turbines that proceeds from Glauert’s theory, but replaces LMADT with a two-dimensional potential flow model following Taylor’s [32] theory for drag on a porous plate, with adjustments for mass conservation [33] and wake pressure [30]. Steiros et al. [24] demonstrated that this approach overcame the limitations of LMADT for higher thrust coefficients, while maintaining good agreement with the more conventional method at lower confinements and solidities. However, when validating their model with drag measurements on porous plates, Steiros et al. [24] observed increasing disagreement between their theory and experiments with increasing confinement and solidity, which leaves open the question of whether a bluff-body correction could be more accurate in this regime.
Experimental Numerical
Author Turbine Type $\beta$ tested [%] Author Turbine Type $\beta$ tested [%]
Takamatsu et al. [34] CFT 75, 95 Nishino and Willden [10] Actuator disk 3.1 - 50.3
Whelan et al. [2] AFT 5, 64 Consul et al. [35] CFT 12.5, 25, 50
McAdam et al. [36, 37] CFT 47, 59 Goude and Ågren [18] CFT 0, 12.5, 25, 50
Battisti et al. [9] CFT 2.8, 10 Kolekar and Banerjee [38] AFT 4.2, 8.2, 11, 17, 20, 33, 42
Chen and Liou [20] AFT 10.2, 20.2, 28.3 Schluntz and Willden [13] AFT 0, 7.9, 19.6, 31.4
Ross and Altman [21] CFT (Savonius) 2, 3.5, 8 Gauthier et al. [16] Oscillating foil 0.2 - 60.5
Birjandi et al. [39] CFT 24.5 - 49.2 Sarlak et al. [14] AFT 2, 5, 9, 20
McTavish et al. [11] AFT 6.3, 9.9, 14.3, 19.4, 25.4 Kinsey and Dumas [17] AFT 0, 10, 20, 50, 60
Gaurier et al. [40] AFT 1.2, 3.3, 4.8 CFT 0, 13, 26, 51
Ryi et al. [41] AFT 8.1, 18.0, 48.1 Badshah et al. [42] AFT 2 - 19
Dossena et al. [12] CFT 2.8, 10 Gauvin-Tremblay and Dumas [43] CFT 5, 20, 40
Jeong et al. [22] CFT 3.5, 13.4, 24.7 Zilic de Arcos et al. [25] AFT 1, 5, 10, 20, 40
Ross and Polagye [23] AFT 2, 35 Abutunis and Menta [19] AFT 4.2, 20.8, 41.6, 62.5
CFT 3, 36 Steiros et al. [24] AFT 3.1, 19.6
Ross and Polagye [15] CFT 14, 36 Zhang et al. [26] AFT 2, 2.8, 4.4, 7.9, 17.7
CFT (Savonius) 2.5, 3.6, 5.6, 10, 22.5
  • Value(s) estimated from provided turbine and channel dimensions.
Table 1: Summary of experimental and numerical studies that consider the effects of varying the blockage ratio for different types of turbines (axial-flow turbines are indicated by “AFT” and cross-flow turbines are indicated by “CFT”). Studies that consider $\beta\geq 30\%$ are bolded.
Although the studies in Table 1 may appear to comprehensively explore the effects of confinement on turbine performance, the upper end of blockage ratios that are achievable in a realistic river or tidal channel ($\beta=30\%-60\%$) have been primarily explored through numerical simulation. While the general trends in turbine performance and wake characteristics observed in these simulations are in agreement with theory, the computational methods employed are often validated only at lower blockage [35, 38, 16, 17]—or not at all—before investigating higher blockages. This is likely due to a lack of equivalent experimental data at high blockage ratios; only a handful of the experimental studies have considered blockage ratios greater than $30\%$. Experimental studies that do explore the upper-end of blockage ratios consider only a sparse set of $\beta$, likely due to difficulty of maintaining dynamic similarity across a wider range of blockage ratios [44]. Furthermore, meaningful cross-comparisons between experimental studies are complicated by differences in rotor geometry, and several studies simultaneously vary blockage along with Reynolds and Froude numbers [e.g., 36, 37, 39]. Consequently, although the general effects of blockage on turbine performance are well-known, there remains uncertainty as to how these effects evolve at higher confinement. Additionally, despite the widespread application of analytical blockage corrections in prior work, their suitability for higher confinements has not been fully established, nor has their ability to describe the flow field around confined turbines been thorough explored.
[…]
Figure 1: Outline of the procedure in Sections III and IV to interpret high blockage cross-flow turbine dynamics using 1 experimental performance data, 1 flow fields, and 1 linear momentum theory, resulting in 1 self-similar array performance across blockage ratios. Key methodologies that are adapted from prior work are highlighted in gray boxes.
The objective of this work is to develop a more comprehensive understanding of performance and flow fields around a cross-flow turbine array at higher blockages than previously considered in most experiments. Such information can be used to understand the relevance of theory to non-ideal arrays that arguably violate multiple assumptions underpinning analytical models (e.g., treating the turbines as porous, non-rotating disks in LMADT). To this end, a symmetric two-turbine array is tested in a recirculating water channel at blockage ratios ranging from $30\%$ to $55\%$, and flow fields in the near wake are characterized using particle image velocimetry (PIV). The Reynolds and Froude numbers are controlled to avoid convolving changes in blockage with other relevant non-dimensional parameters [45]. Section II describes the experimental methods used to characterize array performance and flow fields, with resulting performance, near-wake velocities, and free surface deformation discussed in Section III. In Section IV, the open-channel LMADT model developed by Houlsby et al. [1] is applied to the experimental performance data to model the wake velocity, bypass velocity, and free surface deformation in the vicinity of the array. The analytically-modeled velocities are compared to measurements and are used to interpret high-blockage turbine hydrodynamics following Whelan et al.’s [2] adaptation of the bluff-body theory of Maskell [3]. The procedure followed in Sections III and IV is graphically outlined in Figure 1 and methodologies that are adapted from prior work are highlighted. Based on the characteristic dynamics observed, a method for forecasting confined turbine performance based on LMADT and Maskell’s [3] bluff body theory (i.e., the reverse of the procedure in Figure 1) is presented in Section V.
II Experimental Methods
II.1 Cross-Flow Turbine Test Setup
Experiments are conducted in the Alice C. Tyler recirculating water flume at the University of Washington. The flume has a test section that is 4.88 m long with a width ($w$) of 0.76 m, and can accommodate water depths up to 0.60 m and flow speeds up to ${\sim}\!1.1$ m/s. The flume is equipped with both a heater and chiller for temperature control, and can maintain water temperatures between $10$ ${}^{\circ}\mathrm{C}$ and $40$ ${}^{\circ}\mathrm{C}$.
The laboratory-scale array consists of two identical straight-bladed cross-flow turbines. Experiments with a pair of turbines, rather than a single turbine, are motivated by the research application to confinement benefits for turbine arrays. The rotors are two-bladed, with each blade consisting of a NACA 0018 profile with a 0.0742 m chord length ($c$) mounted at a $-6^{\circ}$ (i.e., toe-out) preset pitch angle as referenced from the quarter chord. The blade span ($H$) is 0.215 m and the turbine diameter ($D$) is 0.315 m, measured at the outermost point swept by the blades. The blades are attached to the central driveshaft of each rotor using thin, hydrodynamic blade-end struts (NACA 0008 profile, 0.0742 m chord length). This arrangement results in an aspect ratio ($H/D$) of 0.68, chord-to-radius ratio ($c/R$) of 0.47, and solidity ($Nc/\pi D$, where $N$ is the blade count) of 0.15.
Figure 2: A rendering of the experimental test-rig, as viewed from upstream. The configuration shown corresponds to a blockage ratio of $40.1\%$.
Figure 3: Overhead view of the array layout in the Tyler flume, with key measured quantities annotated. The outermost circle swept by each turbine (with diameter $D$) is shown as a dashed blue circle.
The two rotors, designated as “Turbine A” and “Turbine B”, are integrated into the experimental setup shown in Figure 3, which consists of two identical test-rigs. The top of each turbine’s central shaft is connected by a flexible shaft coupling (Zero-Max SC040R) to a servomotor (Yaskawa SGMCS-05BC341) which regulates the rotation rate of the turbine. This control strategy yields similar time-average performance [46] and wake characteristics [47] to applying a constant oppositional torque and reduces cycle-to-cycle variability . The angular position of each turbine is measured via the servomotor encoder ($2^{16}$ counts per revolution), from which the angular velocity is estimated. The bottom of each turbine’s central shaft sits in a bearing. The net forces and torques on each turbine are measured by a pair of 6-axis load cells: an upper load cell (ATI Mini45-IP65) connected to the servomotor through a thermal isolation assembly and fixed to a rigid crossbeam, and a lower load cell (ATI Mini45-IP68) mounted to the bottom bearing and fixed to the bottom of the flume via a suction plate. Measurements from the load cells and servomotor encoders for both turbines are acquired synchronously at 1000 Hz in MATLAB using two National Instruments PCIe-6353 DAQs.
The freestream velocity, $U_{\infty}$, is measured using an acoustic Doppler velocimeter (Nortek Vectrino Profiler) sampling at 16 Hz. The velocimeter samples a single cell positioned laterally in the center of the flume, vertically at the turbine midspan, and $5D$ upstream of the turbines’ axes of rotation. Velocity measurements are despiked using the method of Goring and Nikora [48]. The turbulence intensity of the freestream flow measured by the ADV is ${\sim}2\%$ for all experiments. The water depth far upstream of the array ($5.8D$ upstream of the turbines’ axes of rotation, centered laterally in the channel) and depths in vicinity of the array ($0.5D$, $1.0D$, and $2.0D$ directly upstream and downstream of Turbine B) are measured using seven synchronized ultrasonic free surface transducers (Omega LVU32) sampling at 0.5 Hz. As the array is symmetric about the channel centerline (Figure 3) the streamwise free surface profile in the vicinity of Turbine B is assumed to be representative of that in the vicinity of Turbine A. The water temperature is measured using a temperature probe (Omega Ultra-Precise RTD) and maintained within ±$\pm 0.1^{\circ}\mathrm{C}$ of the target value during each experiment.
II.2 Non-Dimensional Parameters
For the two-turbine array, the blockage ratio is defined as:
$\beta=\frac{A_{\mathrm{turbines}}}{A_{\mathrm{channel}}}=\frac{2HD}{hw},$ (2)
where $h$ is the dynamic water depth and $w$ is the channel width. As subsequently discussed, because of the array layout, global and local blockage for this array are identical and can be described by a single value. The blockage ratio is varied by holding the array geometry constant and changing the water depth in the flume. To isolate blockage effects, other non-dimensional flow parameters must be controlled as the depth changes. For example, decreasing water depth (increasing $\beta$) increases the depth-based Froude number,
$Fr_{h}=\frac{U_{\infty}}{\sqrt{gh}}\ \ ,$ (3)
where $g$ is the acceleration due to gravity. The array’s proximity to the free surface also changes as depth decreases, represented here by the normalized submergence, $s/h$, where $s$ is the distance between the free surface and the top of the turbine blades. Both $Fr_{h}$ [35, 49, 45] and $s/h$ [39, 38, 50, 45] have been shown to influence turbine performance. Therefore, as $h$ decreases, $U_{\infty}$ and $s$ must also decrease to hold $Fr_{h}$ and $s/h$, respectively, constant.
However, a decrease in $U_{\infty}$ also decreases the Reynolds number, which is defined here with respect to the freestream velocity and turbine diameter as
$Re_{D}=\frac{U_{\infty}D}{\nu}\ \ ,$ (4)
where $\nu$ is the kinematic viscosity. The dependence of turbine performance on the Reynolds number is well documented [51, 52, 53, 45, 54]. Although turbine performance becomes independent of the Reynolds number above a certain threshold, for cross-flow turbines, this is difficult to achieve at laboratory scale without a compressed-air wind tunnel [51, 52] or relatively large tow tank [53]. To compensate for the decrease in $U_{\infty}$, the kinematic viscosity is decreased by changing the water temperature. In this way, $\beta$ can be varied while holding $Fr_{h}$, $s/h$, and $Re_{D}$ constant.
Target Blockage Condition Flume Parameters Non-Dimensional Flow Parameters
$\beta$ [%] $h$ [m] $U_{\infty}$ [m/s] $s$ [m] Temp [${}^{\circ}\mathrm{C}$] $s/h$ $Fr_{h}$ $Re_{D}$
30.0 0.593 0.528 0.326 21.0 0.55 0.22 $1.70\!\times\!10^{5}$
33.4 0.534 0.501 0.267 23.3 0.50
36.7 0.485 0.477 0.218 25.4 0.45
40.1 0.445 0.457 0.178 27.3 0.40
45.0 0.396 0.431 0.129 30.1 0.33
50.0 0.356 0.409 0.090 32.6 0.25
55.0 0.324 0.390 0.057 35.0 0.18
Table 2: Experimental parameters for each blockage condition tested.
Array performance is measured at blockage ratios from $30.0\%$ to $55.0\%$, which is the widest range possible given the size of these turbines and flume capabilities. The inflow conditions required to achieve each blockage while holding $Fr_{h}$ and $Re_{D}$ constant are summarized in Table 2. Across all experiments, the measured $\beta$ are within $1.5\%$ of the target values in Table 2, and the measured $Fr_{h}$ and $Re_{D}$ do not deviate more than $5\%$ from the target values.
The value of $Re_{D}$ in this study is constrained by the maximum $U_{\infty}$ for the highest blockage ratio. This maximum velocity, in turn, is constrained by the entrainment of air from the free surface into the rotor (i.e., ventilation), which decreases lift on the blades and degrades performance [39, 55]. Consequently, the maximum $U_{\infty}$ for the highest tested blockage ratio is set such that any ventilation occurs beyond the rotation rate corresponding to maximum turbine performance. To further limit the risk of ventilation, $s/h$ is maximized at each blockage rather than held constant across all blockages, since ventilation becomes more likely as $s/h$ decreases. Array performance at various $s/h$ and $\beta$ is separately evaluated (Appendix A) and confirms that varying $s/h$ has minor effects on performance at these test conditions before the onset of ventilation.
II.3 Array Layout and Control
An overhead view of the array layout is shown in Figure 3. The turbines are positioned laterally such that the center-to-center spacing is ${\sim}1.2D$ and the blade-to-blade spacing between adjacent turbines (${\sim}0.2D$; ${\sim}0.9c$) is twice the wall-to-blade spacing (i.e., the walls notionally correspond to symmetry planes in a larger array). The turbines are operated under a counter-rotating, phase-locked scheme, wherein both turbines rotate at the same, constant speed, but in opposite directions (i.e., $\omega_{A}=-\omega_{B}$), with zero phase offset between them. The turbines are counter-rotated such that the blades rotate toward the channel centerline, which has been shown to augment performance relative to other rotation schemes [56, 57, 58].
II.4 Performance Metrics
Array performance metrics are calculated from the measured quantities shown in Figure 3. The rotation rate is non-dimensionalized as the ratio of the blade tangential velocity to the freestream velocity, or the tip-speed ratio, as
$\lambda=\frac{\omega R}{U_{\infty}}\ \ ,$ (5)
where $\omega=|\omega_{A}|=|\omega_{B}|$ is the magnitude of the angular velocity. Data are collected at each tip-speed ratio for 60 seconds, and the time series is cropped to an integer number of turbine rotations.
The array efficiency (power coefficient) is the ratio of net mechanical power produced by the turbines to the kinetic power in the freestream flow that passes through array’s projected area
$C_{P}=\frac{Q_{A}\omega_{A}+Q_{B}\omega_{B}}{\frac{1}{2}\rho U_{\infty}^{3}A_{\mathrm{turbines}}}\ \ ,$ (6)
where $Q_{A}$ and $Q_{B}$ are the the hydrodynamic torques on the turbines and $\rho$ is the density of the working fluid. Structural loads on the array are characterized via the thrust coefficient
$C_{T}=\frac{T_{A}+T_{B}}{\frac{1}{2}\rho U_{\infty}^{2}A_{\mathrm{turbines}}}\ \ ,$ (7)
where $T_{A}$ and $T_{B}$ are the streamwise forces on the turbines. Cross-flow turbines also experience forces in the cross-stream direction ($L_{A}$ and $L_{B}$ in Figure 3), which can be similarly non-dimensionalized as an array lateral force coefficient using an analogous equation to Eq. 7. The lateral forces are tangential to this work, but provided as supplementary material [59].
II.5 Flow Visualization
Figure 4: Schematic of the PIV test setup including turbines and example flow field data.
Figure 5: Location of PIV field of view within the flume, looking downstream. The lateral and vertical coordinates for the field of view are defined with respect to the center of Turbine B. The black rectangles indicate the projected areas of the turbines. The purple lines indicate the nominal location of the free surface at each blockage ratio.
Two-camera, three-component particle image velocimetry is used to measure the near-wake flow field in planes perpendicular to the flow (Figure 5). Two cameras (Imager sCMOS 5.5 MPx, maximum 50 fps) with 35 mm and 60 mm lenses, are positioned on either side of the flume at approximately $45^{\circ}$ incidence to the image plane. The cameras are oriented perpendicular to water-filled acrylic prisms shown in Figure 5, which are used to reduce refractive distortion from the angled camera positioning relative to the flume. Scheimpflug camera lens adaptors correct any residual focal distortions from angled fields of view (FoV). The FoV of each camera is focused on a domain centered behind one of the turbines and results in an overlapping 42.3$\times$27.7 cm2 data collection region. The resulting FoV (Figure 5) is positioned approximately 5.2 cm above the bottom of the flume and approximately 8.75 cm from the side wall to maximize capture of the flow through and around the array while reducing overexposure from laser scattering by solid boundaries. The FoV encompasses the majority of the projected area of Turbine B, the bypass region above Turbine B, and the bypass region between the turbines. Although the FoV does not capture the full array, given the layout symmetry, the flow measured around Turbine B is expected to be descriptive of that near Turbine A. The entire flume is seeded with neutrally buoyant hollow glass spheres with a diameter of $\sim\!10\mu\mathrm{m}$ (Potters Industries Sphericel 110P8), and illumination is provided by a 532 nm Nd:YAG laser (EverGreen 200), which produces a light sheet approximately 2 mm thick.
The near-wake flow field is measured in vertical planes $0.6D$ and $1.5D$ downstream of the turbines’ axes of rotation, with the array moved in the streamwise direction to change the relative location of the vertical plane. We focus our analysis on the plane nearest to the array at $0.6D$, with the flow fields at $1.5D$ provided in supplementary material [59]. Flow field data in each plane are collected at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ for tip-speed ratios of $\lambda=1.6$, $2.6$, $2.9$, and $3.6$ (all $\beta$) and $\lambda=4.2$ ($\beta=55.0\%$ only). These tip-speed ratios include the “optimal” $\lambda$ corresponding to maximum $C_{P}$ for $\beta=36.7\%$, $45.0\%$, and $55.0\%$, as well as $\lambda$ lower (“underspeed”) and higher (“overspeed”) than the optimal $\lambda$ at these blockage ratios, which are relevant in field deployments for regulating turbine power above the rated flow speed [60]. Data acquisition is synchronized with turbine phase ($\theta$), with images collected in $3^{\circ}$ increments at the $\lambda$ corresponding to maximum $C_{P}$ for $\beta=36.7\%$, $45.0\%$, and $55.0\%$, and $12^{\circ}$ increments for all other $\lambda$ cases. A total of 40 frames are collected at each phase and are averaged to reduce measurement noise. All phases are then averaged together to produce representative time-averaged fields.
LaVision Davis 10.2.1 software is used to acquire and process the PIV data. A dual-plane calibration target and self-calibration are employed. The resulting scale factor is 7.97 pix/mm with fitting errors of 0.81 and 0.92 pixels for each camera, respectively. A minimum filter subtracts background noise and an average polynomial filter using an interrogation window of 11$\times$11 pixels minimizes unsteady reflections from the free surface or near-blade regions. Vector calculation employs an initial $128\times 128$ interrogation window with a 50% overlap followed by four sequential reducing passes to $32\times 32$ pixels at a 75% overlap. After vector calculation, three passes of a universal outlier detection median filter [61] removes vectors with a median normalized residual greater than 1.5 on a filter domain of $7\times 7$ vectors. These vectors are reinstated if their normalized residual is less than 2.5 and alternate correlation peaks have a higher residual. Groups with fewer than five vectors are removed.
III Experimental Results
III.1 Array Performance
[…]
Figure 6: Time-averaged 6 $C_{P}$ and 6 $C_{T}$ as a function of $\beta$ and $\lambda$. The shaded regions indicate the interquartile range of the array- and cycle-averaged performance at each $\beta$ and $\lambda$ (the vertical span of the shaded region at each point is similar to the size of the markers for most operating conditions).
Time-average $C_{P}$ as a function of $\beta$ and $\lambda$ is shown in Figure 6. As for prior studies (Table 1), as $\beta$ increases, the maximum $C_{P}$ increases, the array produces power over a broader range of tip-speed ratios, and the optimal tip-speed ratio increases. Similarly, the time-averaged $C_{T}$ (Figure 6) generally increases with $\beta$, particularly at higher blockage ratios. Above $\beta=33.4\%$, maximum $C_{P}$ exceeds the Betz limit, and above $\beta=50.0\%$, maximum $C_{P}$ exceeds unity. Such efficiencies are not violations of energy conservation since the definition of $C_{P}$ in Eq. 6 considers only the kinetic power passing through the array projected area. This neglects the fluid’s potential energy, which is appreciably drawn down as $\beta$ and $C_{T}$ increase. As $\beta$ increases, a unity-bounded efficiency metric may resemble the hydraulic efficiency of a hydropower turbine (e.g., metrics employed by Takamatsu et al. [34] or McAdam et al. [36]) in which the available power is a function of volumetric flow rate through the channel and the head drop across the array. This alternative efficiency metric is explored further in the supplementary material [59], but is sensitive to the resolution and accuracy of free surface measurements. Here, the conventional definitions of $C_{P}$ and $C_{T}$ facilitate comparison with prior studies, as well as consistency with LMADT—which is formulated using the definition of $C_{T}$ in Eq. 7.
[…]
Figure 7: Time-averaged 7 $C_{P}$ and 7 $C_{T}$ versus $\beta$ at constant $\lambda$. The data for each $\beta$ are interpolated at each $\lambda$ from the points in Figure 6. The dashed lines represent the line of best fit through the data points corresponding to $\beta\leq 45.0\%$ at each $\lambda$.
When the array power and thrust coefficients are regressed against $\beta$ at constant tip-speed ratio, approximately linear relationships are revealed, which are shown for a subset of $\lambda$ in Figure 7. Below a threshold value of $\lambda$, both $C_{P}$ and $C_{T}$ are independent of the blockage ratio and depend only on the tip-speed ratio, which corresponds to the collapsed regions of the $C_{P}-\lambda$ curves ($\lambda<1.6$) and $C_{T}-\lambda$ curves ($\lambda<1.2$) in Figure 6. Similar relations are observed in prior work for both cross-flow turbines[35] and axial-flow turbines [38, 42]. As $\lambda$ increases further, $C_{P}$ and $C_{T}$ vary linearly with the blockage ratio for $\beta\leq 45.0\%$ and moderate $\lambda$, which is emphasized by the linear regressions in Figure 7. While similar linear relationships between the blockage ratio, power, and thrust have been previously identified in simulations of oscillating foils [16], axial-flow turbines [17, 19] and cross-flow turbines [17], this is the first experimental demonstration of this trend. Finally, as $\beta$ and $\lambda$ increase further, these relationships become non-linear with performance increasingly sensitive to incremental changes in the blockage and tip-speed ratios, which is similar to narrower observations for $C_{P}$ by Gauthier et al. [16].
III.2 Flow Fields
[…]
Figure 8: Time-averaged streamwise velocity $0.6D$ downstream of the array as function of $\beta$ and $\lambda$. Velocity measurements are normalized by the freestream velocity measured by the ADV. The black rectangles in each tile indicate the projected area of the turbines in the field of view. Flow fields corresponding to the optimal $\lambda$ for each blockage are outlined in purple (tiles 8 8, and 8).
Figure 8 shows the normalized time-average streamwise velocity field $0.6D$ downstream of the array (i.e., within the near-wake) as a function of $\beta$ and $\lambda$. The velocity fields are characterized by a wake region behind the turbines ($U_{x}/U_{\infty}<1$, indicated by blue hues) and a bypass region with accelerated flow above, below, and between the turbines ($U_{x}/U_{\infty}>1$, indicated by red hues). The bypass flow region between Turbine B and the flume wall lies outside of the field of view, and we attribute the persistent region of decelerated flow below Turbine B at all $\lambda$ to the test rig wake (Figure 3).
[…]
Figure 9: Time-averaged normalized streamwise velocity profiles measured $0.6D$ downstream of the array for various $\beta$ at 9 $\lambda=1.6$, 9 $\lambda=2.6$, and 9 $\lambda=3.6$. The profiles correspond to the average velocity along the middle $70\%$ of the turbine blade span. Vertical gray rectangles correspond to the bypass region.
Trends in the evolution of the near-wake with $\beta$ and $\lambda$ are apparent in two-dimensional velocity profiles of the average streamwise velocity at each $y/D$ (Figure 9). These profiles are calculated over the middle $70\%$ of the blade span to exclude edge effects from the blade-end struts. Ripples in the wake region of these profiles are attributed to the passage of vortices shed from the blades during their downstream sweep. At $\lambda=1.6$ (Figure 9), where array thrust varies weakly with $\beta$ (Figure 7), $U_{x}/U_{\infty}$ increases slightly with $\beta$ across the lateral profile. As $\lambda$ increases (Figures 9 and 9), the turbines impart greater resistance on the flow and more fluid is diverted around them, causing $U_{x}/U_{\infty}$ to increase in the bypass region for all $\beta$. Similarly, at a given $\lambda$, the bypass velocity increases with $\beta$. $U_{x}/U_{\infty}$ in the wake region also tends to increase with $\beta$ at a given $\lambda$ since confinement increases the flow through the turbines. However, increases in the wake velocity with $\beta$ are less pronounced than the corresponding increases in bypass velocity due to momentum extraction by the array from the core flow through the turbines. Although the maximum bypass velocity at the array centerline increases with both the $\beta$ and $\lambda$, $U_{x}/U_{\infty}$ decreases more rapidly away from this midpoint at higher $\lambda$. This is indicated by the narrower peaks in the velocity profiles as $\lambda$ increases at a given $\beta$. Conversely, an increase in $\beta$ at constant $\lambda$ results in a more gradual reduction of $U_{x}/U_{\infty}$ away from the array centerline, which implies increased turbulent mixing between the bypass flow and flow through the turbine. As discussed by Nishino and Willden [10], mixing between the bypass and core flows provides a secondary power enhancement with increasing blockage by increasing the pressure drop across the rotor as a consequence of elevated near-wake velocity.
III.3 Free Surface Deformation
[…]
Figure 10: Time-averaged streamwise water depth profiles for various $\lambda$ at 10 $\beta=36.7\%$, 10 $\beta=45.0\%$, and 10 $\beta=55.0\%$. The farthest upstream measurement is taken at the center of the channel, whereas depth measurements in the vicinity of the array are directly upstream and downstream of the Turbine B (i.e., $y/D=0$ in the PIV flow fields). Depth measurements at each streamwise location ($h_{x}$) are normalized by the far-upstream water depth ($h$) measured at the same $\beta$ and $\lambda$. The dashed lines indicate the extent of the turbine’s swept area.
The time-average streamwise free surface profile across Turbine B is shown in Figure 10 at the same $\beta$ and $\lambda$ as for the flow fields in Figure 8. For all $\beta$ and $\lambda$, a free surface drop is observed across the turbine, corresponding to the pressure drop across a turbine in an open-channel flow [6]. The magnitude of this drop (in this case, the difference between the measured water depths at $x/D=-1$ and $x/D=1$) tends to increase with $\beta$ for a given $\lambda$. However, as shown in Appendix A, the free surface drop behind the array also depends on the normalized submergence depth ($s/h$) which decreases with $\beta$ in these experiments (Section II.2). Consequently, the free surface drop at higher $\beta$ is likely augmented by the closer proximity of the turbines to the free surface. The free surface drop across the turbine clearly increases with $\lambda$ at $\beta=55.0\%$ (Figure 10), consistent with the monotonic increase in $C_{T}$ with $\lambda$ (Figure 6). However, even though similar relationships between $C_{T}$ and $\lambda$ are observed for $\beta=36.7\%$ and $\beta=45.0\%$, the free surface drop at these blockages does not vary significantly with the tip-speed ratio for $\lambda\geq 2.6$, although the depth at $x/D=0.5$ does continuously decrease with $\lambda$ for all $\beta$.
IV Evaluation of LMADT for a Cross-flow Turbine Array at High Blockage
IV.1 Model Overview
Figure 11: Linear momentum model for an actuator disk in open-channel flow defined by Houlsby et al. [1] and adapted to the cross-flow turbine array nomenclature.
The performance, flow field, and free surface measurements described in Section III allow us to directly assess the accuracy of a one-dimensional linear momentum actuator disk theory (LMADT) description of the cross-flow turbine array. Given the appreciable free surface deformation (Figure 10) and the homogeneous lateral distribution of turbines in the channel (i.e., equivalent local and global blockage), the open-channel LMADT model introduced by Houlsby et al. [1] is most relevant. In the limiting case of negligible free surface drop across the turbine array ($Fr_{h}^{2}\rightarrow 0$), Houlsby et al.’s model reduces to the well-known model of Garrett and Cummins [4]. For a non-uniform distribution of turbines along the channel width, a theory that represents the individual turbines in the array as separate actuator disks would be required [e.g., 5, 6, 8]. Although the array’s homogeneity does not extend to the vertical direction, the blockage asymmetry in these experiments (1.3 - 2.3) is within the range found by Kinsey and Dumas [17] to have limited influence on turbine power and thrust. We note that LMADT-based models like that of Houlsby et al. [1] are only one of several types of analytical models (e.g., double-multiple streamtube models [62, 63], potential flow models [24]) that could be applied. The potential flow model of Steiros et al. [24] is given separate consideration in supplementary material [59].
In the LMADT model, the array is represented as a single actuator disk with the same blockage ratio and thrust coefficient as the array (Figure 11). The flow upstream of the turbine is unidirectional, uniform, and subcritical (i.e., $Fr_{h}^{2}<1$) with velocity $U_{\infty}$ and depth $h$. In response to momentum extraction by the turbine, the “core” flow in the streamtube that passes through the rotor decelerates from $U_{\infty}$ to $u_{t}$ just upstream of the rotor plane, and further decelerates to $u_{w}$ downstream of the turbine in the near-wake. To satisfy continuity, the fluid outside of this streamtube is accelerated from $U_{\infty}$ to $u_{b}$. The core wake velocity ($u_{w}$) and bypass velocity ($u_{b}$) are both defined at a location downstream of the turbine where the pressure in the core and bypass flows is equal and hydrostatic. The flow is assumed to be axisymmetric and inviscid up to this location, downstream of which $u_{w}$ and $u_{b}$ mix. Far downstream, the flow is once again uniform, with reduced depth due to the energy extracted by the disk but higher velocity to conserve mass. This model qualitatively captures aspects of the experimentally observed wake and bypass velocities (Figure 8). However, although LMADT models are widely used to represent turbine hydrodynamics in confined flows, several of the underlying assumptions are often violated. For example, the wake is assumed to be axisymmetric and to not mix with the bypass until static pressure has equilibrated between the core and bypass flows, while we observe a non-uniform wake with immediate mixing between core and bypass.
Detailed derivations of the LMADT equation set used here are presented in Houlsby et al. [1] and Houlsby and Vogel [64]. The implementation in this study follows Ross and Polagye [23], who reorganized Houlsby et al.’s original formulation into two equations from which $u_{w}$ and $u_{b}$ can be obtained numerically if $U_{\infty}$, $C_{T}$, $Fr_{h}$, and $\beta$ are known:
$u_{w}=\frac{Fr_{h}^{2}u_{b}^{4}-(4+2Fr_{h}^{2})U_{\infty}^{2}u_{b}^{2}+8U_{\infty}^{3}u_{b}-4U_{\infty}^{4}+4\beta C_{T}U_{\infty}^{4}+Fr_{h}^{2}U_{\infty}^{4}}{-4Fr_{h}^{2}u_{b}^{3}+(4Fr_{h}^{2}+8)U_{\infty}^{2}u_{b}-8U_{\infty}^{3}}\ ,$ (8)
$u_{w}=\sqrt{u_{b}^{2}-C_{T}U_{\infty}^{2}}\ .$ (9)
The velocity through the turbine, $u_{t}$, is then given as
$u_{t}=\frac{u_{w}(u_{b}-U_{\infty})(2gh-u_{b}^{2}-u_{b}U_{\infty})}{2\beta gh(u_{b}-u_{w})}\ .$ (10)
For the solutions to Eqs. 8, 9 and 10 to be physically valid, the core and bypass flows must be subcritical and satisfy $u_{b}>u_{t}>u_{w}$ and $u_{b}>U_{\infty}>u_{t}$.
The free surface profile in the vicinity of the array can also be calculated from the open-channel LMADT model. Houlsby et al. define five streamwise stations, $x_{1}-x_{5}$:
  1. $\mathbf{x_{1}}$
    Far upstream, where the flow is undisturbed by the disk ($h_{1}$ = $h$; $U_{1}=U_{\infty}$);
  2. $\mathbf{x_{2}}$
    Just upstream of the actuator disk ($h_{2}$; $U_{2}=u_{t}$);
  3. $\mathbf{x_{3}}$
    Just downstream of the actuator disk ($h_{3}$; $U_{3}=u_{t}$);
  4. $\mathbf{x_{4}}$
    Downstream of the actuator disk, where the static pressure in the core and bypass
  5. flows is equal ($h_{4}$; $U_{4}$ = $u_{w}$ in the core flow; $U_{4}=u_{b}$ in the bypass); and
  6. $\mathbf{x_{5}}$
    Far downstream of the disk, when the flow has fully mixed ($h_{5}$; $U_{5}$),
where $h_{1}$ through $h_{5}$ and $U_{1}$ through $U_{5}$ are the water depth and velocity, respectively, at each station. Once $u_{w}$, $u_{b}$, and $u_{t}$ have been calculated from Eqs. 8, 9 and 10, then $h_{4}$, $h_{3}$, and $h_{2}$ may be calculated following Houlsby et al. [1] as
$h_{4}=h+\frac{(U_{\infty}^{2}-u_{b}^{2})}{2g}\ \ ,$ (11)
$h_{3}=h_{4}+\frac{(u_{w}^{2}-u_{t}^{2})}{2g}\ \ ,$ (12)
$h_{2}=h+\frac{(U_{\infty}^{2}-u_{t}^{2})}{2g}\ \ .$ (13)
The normalized net free surface drop across the channel from far upstream ($x_{1}$) to far downstream ($x_{5}$), $\frac{\Delta h}{h}$, is obtained by solving
$\frac{1}{2}\left(\frac{\Delta h}{h}\right)^{3}-\frac{3}{2}\left(\frac{\Delta h}{h}\right)^{2}+\left(1-Fr_{h}^{2}+\frac{C_{T}\beta Fr_{h}^{2}}{2}\right)\frac{\Delta h}{h}-\frac{C_{T}\beta Fr_{h}^{2}}{2}=0\ \ ,$ (14)
from which $h_{5}$ is obtained as
$h_{5}=h\left(\frac{\Delta h}{h}\right)\ \ .$ (15)
In our experiments, because the static pressure equilibrium point is unknown, the location of $x_{4}$ is ambiguous, as are the locations of $x_{2}$ and $x_{3}$ due to the non-zero rotor “thickness”. In addition, the flume has insufficient length for the flow to fully mix downstream of the array, such that $x_{5}$ cannot be measured. Consequently, although $u_{b}$, $u_{w}$, and $u_{t}$ may be estimated from available measurements and $h_{2}$ through $h_{5}$ calculated, the corresponding streamwise positions of $x_{2}$ through $x_{5}$ in experiments are uncertain.
IV.2 Comparison between Experimental and Modeled Near-Wake Velocities
[…]
Figure 12: The 12 bypass velocities, and 12 wake velocities (all normalized by the freestream velocity) predicted by the open-channel LMADT model of Houlsby et al. [1] from measurements of array thrust and flow conditions.
Using the measured values of $C_{T}$, $U_{\infty}$, $Fr_{h}$, and $\beta$, $u_{b}$ and $u_{w}$ are calculated at each $\lambda$ using Eqs. 8 and 9. The resulting analytically-predicted velocities are normalized by the freestream velocity in Figure 12. At a given tip-speed ratio, both $u_{b}$ (Figure 12) and $u_{w}$ (Figure 12) increase with $\beta$ because confinement accelerates the flow through and around the turbine. $u_{b}$ increases with $\lambda$ at all $\beta$ since the corresponding increase in thrust diverts more flow around the array, whereas $u_{w}$ decreases with $\lambda$ at all $\beta$ due to momentum extraction by the array. In agreement with experimental trends in the velocity profiles in Figure 9, modeled $u_{b}$ increases more strongly with tip-speed ratio at higher blockage. We note that the modeled $u_{w}$ is always positive, indicating that the upper limit of the thrust coefficient for physically-valid solutions to open-channel LMADT (analogous to $C_{T}\leq 1$ for LMADT in unconfined flow) is not exceeded in these experiments.
[…]
Figure 13: Time-averaged streamwise velocities measured $0.6D$ downstream of the array at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ and optimal $\lambda$, normalized by the corresponding 13-13 modeled bypass velocity, $u_{b}$, and 13-13 modeled wake velocity, $u_{w}$. The black rectangles indicate the projected area of the turbines in the field of view.
[…]
Figure 14: Time-averaged streamwise velocity profiles measured $0.6D$ downstream of the array normalized by the modeled bypass velocity at 14 $\beta=36.7\%$, 14 $45.0\%$, and 14 $55.0\%$ and various $\lambda$. The profiles shown correspond to the average velocity along the middle $70\%$ of the turbine blade span. Vertical gray rectangles correspond to the bypass region.
In Figure 13, the time-averaged near-wake velocity fields at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ and optimal $\lambda$ are normalized by $u_{b}$ and $u_{w}$, respectively. Blue hues correspond to experimental velocities lower than the LMADT-modeled value, red hues correspond to experimental velocities higher than the LMADT-modeled value, and white indicates agreement between the experimental and modeled velocities. The measured bypass velocity is well-predicted by LMADT, as indicated by the substantive white regions in Figures 13, 13 and 13. As shown in the bypass-normalized velocity profiles in Figure 14, the modeled $u_{b}$ tends to be slightly higher than experimental measurements at the array centerline, but are within approximately $8\%$ of the measured value for all $\beta$ and $\lambda$. In contrast, the measured wake velocities are lower than modeled wake velocities at all $\lambda$, and there is no substantive region of the flow that is well represented by $u_{w}$ (Figures 13, 13 and 13). Flow fields normalized by $u_{b}$ and $u_{w}$ at other $\beta-\lambda$ combinations are provided as supplementary material [59]. While it is unknown if the static pressure is equal in the wake and bypass flows at $0.6D$ downstream of the array, $u_{b}$ is less descriptive of the measured bypass flow further downstream ($1.5D$), as shown in supplementary material [59]).
[…]
Figure 15: Examples of non-idealities in the measured flow fields that are not present in the flow fields assumed by linear momentum theory, shown for the streamwise flow through a plane $0.6D$ downstream of the array operating at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ and optimal $\lambda$. 1515 Standard deviation of the measured streamwise velocity field, calculated as the standard deviation of the phase-average streamwise velocity field across all phases. 1515 Time-averaged streamwise vorticity at the same location and conditions, calculated from the phase-averaged velocity fields and averaged across all phases. 1515 Time-averaged turbulent kinetic energy (TKE) measured at the same location and conditions. To account for the inherent periodicity of cross-flow turbine wakes, the TKE field shown is obtained by calculating the TKE at each phase from the individual frames collected at that phase, then averaging the resulting fields across all phases.
The quantitative agreement between the modeled and measured bypass velocities is remarkable given the multiple violations of LMADT in the near-wake flow field. First, unlike the uniform and one-dimensional flow assumed by the LMADT model, the flow in the near-wake varies in space and time, even within the bypass and wake regions (Figures 15, 15 and 15). Second, the flow is rotational (Figures 15, 15 and 15): dynamic stall vortices are shed into the region between the two turbines and strong tip vortices form along the top and bottom of the turbine projected area. The asymmetric vorticity produced by these sources drives flow toward the center of each turbine, contributing to wake asymmetry [65, 66]. Third, as suggested by the velocity profiles in Figure 9 and the turbulent kinetic energy (TKE) in the near-wake (Figures 15, 15 and 15), mixing occurs between the bypass flow and the turbine wake, consistent with the shear layer in a co-flowing turbulent jet. In contrast, the LMADT model assumes that the core and bypass flows do not mix until further downstream (Figure 11). This near-wake mixing necessarily reduces the bypass velocity, which may explain why the modeled bypass velocity is slightly higher than the measured bypass velocity at the array centerline (Figure 14).
IV.3 Comparison between Experimental and Modeled Free Surface Deformation
[…]
Figure 16: Modeled free-surface profile at various stations along the channel predicted using the open-channel LMADT model of Houlsby et al. [1] and the measured $C_{T}$, $h$, and $U_{\infty}$ for 16 $\beta=36.7\%$, 16 $\beta=45.0\%$, and 16 $\beta=55.0\%$ for various $\lambda$. The dashed black line indicates the relative location of the array, represented in the linear momentum model as a thin actuator disk.
For comparison with the measured free surface profiles in Figure 10, the theoretical free surface profiles are shown for the same $\beta$ and $\lambda$ in Figure 16. As noted in Section IV.1, we emphasize that the exact locations of $x_{2}-x_{4}$ relative to the physical array are ambiguous and that $x_{5}$ is likely further downstream than can be measured given the finite flume length. Consequently, the locations of $x_{1}$ through $x_{5}$ relative to the actuator disk model of the array shown in Figure 16 are qualitative. With these caveats, we consider the trends in measured free surface deformation relative to model predictions. For all $\beta$, the modeled water depth just upstream of the actuator disk ($h_{2}/h$) increases continuously with $\lambda$. Similarly, the immediate static head drop across the actuator disk ($(h_{3}-h_{2})/h$) and the total free surface drop along the channel ($\Delta h/h$) increase continuously with both $\lambda$ and $\beta$.
This analytical result is consistent with a direct relationship between the pressure drop across the array (i.e., the thrust coefficient) and a drop in the free surface across the array, but conflicts with measured free surface profiles in Figure 10. Although the measured $C_{T}$ increases continuously with $\lambda$ at all $\beta$, the measured free surface drop across the turbines at $\beta=36.7\%$ (Figure 10) and $45.0\%$ (Figure 10) does not depend significantly on $\lambda$ for $\lambda\geq 2.6$. This may suggest that LMADT is not fully descriptive of the evolution of the free surface in the vicinity of high blockage turbines, although the limited spatial and temporal resolution of the instrumentation employed precludes a more comprehensive characterization of the free surface dynamics in this data set. Experimental trends are consistent with modeled free surface deformation for $\beta=55.0\%$, which suggests that smaller-scale spatial variability could be masking expected trends at lower $\beta$.
IV.4 Bluff-Body Scaling of Performance Data
Given that open-channel LMADT is descriptive of the measured bypass flow across a range of array operating conditions, we turn to the question of whether LMADT also describes the measured array performance. In prior work, LMADT is commonly used in analytical blockage corrections to estimate unconfined performance from measurements or simulations of confined operation [e.g., 67, 20, 17, 23]. However, since a perfect blockage correction would be expected to collapse the $C_{P}$ and $C_{T}$ measured at different blockage ratios (e.g., Figure 6) onto common unconfined $C_{P}-\lambda$ and $C_{T}-\lambda$ curves, such corrections imply not just a connection between confined and unconfined performance, but rather a broader self-similarity in turbine performance across blockage ratios. Here, we explore such self-similar aspects of array performance with a focus on the significance of the bypass flow.
[…]
Figure 17: Time-averaged array performance from Figure 6 scaled by 17 the analytical bypass velocities obtained using the model of Houlsby et al. [1], 17 the unconfined freestream velocities from the model of Houlsby et al. [1], 17 the analytical bypass velocities obtained using the model of Steiros et al. [24], and 17 the unconfined freestream velocities from the model of Steiros et al. [24].
To investigate the relationship between the bypass flow and array performance, we apply a method introduced by Whelan et al. [2] and inspired by the bluff-body theory of Maskell [3] to scale the experimental array performance by the corresponding bypass velocities modeled in LMADT:
${C_{P}}^{b}=C_{P}{\left(\frac{U_{\infty}}{u_{b}}\right)}^{3},$ (16)
${C_{T}}^{b}=C_{T}{\left(\frac{U_{\infty}}{u_{b}}\right)}^{2},$ (17)
${\lambda}^{b}=\lambda\left(\frac{U_{\infty}}{u_{b}}\right).$ (18)
As shown in Figure 17, the resulting ${C_{P}}^{b}-{\lambda}^{b}$ and ${C_{T}}^{b}-{\lambda}^{b}$ curves collapse across the tested blockage and tip-speed ratios, indicating that array power, thrust, and their relationship to rotation rate scale with the bypass velocity for these experimental conditions. Consequently, one may interpret $u_{b}/U_{\infty}$ as a similarity variable for array performance that is well-predicted by LMADT. Some residual variation in the bypass-scaled thrust curves is observed, particularly at lower $\lambda^{b}$ (maximum $\Delta{C_{T}}^{b}\approx 0.06$), but the bypass-scaled power curves exhibit excellent collapse at all $\lambda^{b}$ (maximum $\Delta{C_{P}}^{b}\approx 0.016$). As $u_{b}$ is calculated using the measured $C_{T}$ in Eqs. 8 and 9, it is not surprising that $C_{T}$ scales with $u_{b}$. However, given that measurements of $C_{P}$ are not involved in the calculation of $u_{b}$, the collapse in the ${C_{P}}^{b}-{\lambda}^{b}$ curves is notable, especially since the rotation, torque, and lateral forces experienced by the cross-flow turbines are all neglected in LMADT.
The hydrodynamics that underlie the connection between the bypass velocity and array performance are well-described by Nishino and Willden [10]. To summarize their conclusions, as $\beta$ increases, the velocity of the fluid bypassing the array increases (Figures 9 and 12). Since energy is mostly conserved in the bypass flow, this requires an increased pressure drop in the bypass. To satisfy pressure equilibrium between the core and bypass flows both far upstream and far downstream, the pressure drop in the core flow through the turbines must also increase. In an open-channel flow, this pressure drop manifests as a free surface drop along the channel. As $\beta$ increases, the increased pressure drop across the array accelerates the flow through the turbines, which increases power and thrust at higher blockage.
This connection between the bypass velocity and array performance parallels the theory of Maskell [3], who proposed that the drag force on a bluff body in confined flow is related to the flow bypassing the body, not the undisturbed upstream condition. Assuming a uniform and unidirectional inflow and axisymmetric wake, Maskell developed a model for this behavior by combining conservation of momentum in the flow outside the wake with an empirical relationship for wake contraction due to confinement, and demonstrated through experiments that this model was descriptive of the drag on square flat plates at relatively low blockage ratios ($\beta\leq 10\%$). Maskell leveraged this relationship to propose a blockage correction that predicted the drag on a bluff body in unconfined flow given thrust measurements on the same body in a wind tunnel. As the wake characteristics of wind and water turbines resemble those of a bluff body under certain conditions [68, 69, 70], several subsequent studies have extended Maskell’s theory. While some have directly applied Maskell’s flat plate blockage correction to turbine performance data, [25, 26], others have adapted this theory to specific turbine designs through analogous empirical relationships [71, 21, 22, 26].
Whelan et al. [2] were the first to combine the core idea of Maskell’s theory—that the thrust on a bluff body in confined flow scales with the bypass flow—with LMADT to develop a blockage correction for “highly loaded” turbines. Rather than directly implementing Maskell’s blockage correction, Whelan et al. utilized an open-channel LMADT model to calculate the expected $u_{b}$ from axial-flow turbine thrust measurements at high blockage ($\beta=64\%$), and utilized Eqs. 17 and 18 to scale the measured thrust coefficients and tip-speed ratios by this bypass velocity. The resulting ${C_{T}}^{b}-{\lambda}^{b}$ data agreed closely with the unconfined thrust coefficients predicted for the same turbine simulated using blade element momentum theory. Kinsey and Dumas [17] applied the same correction to CFD-simulated thrust data for a cross-flow turbine at $\beta=13\%$, $25\%$, and $51\%$ with $u_{b}$ calculated via closed-channel LMADT. The resulting ${C_{T}}^{b}-\lambda^{b}$ curves yielded better agreement with the simulated unconfined thrust coefficients than was obtained with the popular Barnsley and Wellicome [28] blockage correction. While Whelan et al. left open the question of whether power could be similarly corrected, Ross and Polagye [23] introduced Eq. 16 as an analogous scaling for $C_{P}$. Using the LMADT model of Houlsby et al. [1] to calculate $u_{b}$, Ross and Polagye scaled the corresponding performance data for both an axial-flow turbine ($\beta=2\%$ and $35\%$) and cross-flow turbine ($\beta=3\%$ and $36\%$). However, in contrast to Whelan et al. and Kinsey and Dumas, at tip-speed ratios beyond optimal ($\lambda>5$ for axial-flow, $\lambda>1.85$ for cross-flow) Ross and Polagye observed moderately poor agreement between the measured low-blockage turbine performance and ${C_{P}}^{b}$ and ${C_{T}}^{b}$ calculated from high-blockage turbine performance. Relative to prior work, the bypass-scaled power and thrust coefficients in Figure 17 exhibit the strongest collapse across blockages observed to date, and highlight similarities between array hydrodynamics and bluff-body dynamics for a wide range of high-blockage operating conditions.
Since the scalings in Eqs. 16, 17 and 18 are typically interpreted as a Maskell-inspired bluff-body blockage correction, it is relevant to compare the bypass-scaled performance (Figure 17) to that obtained via a “standard” LMADT blockage correction based on the theory of Glauert [27], which assumes that the thrust on the turbine responds to the velocity through the turbine ($u_{t}$), not the bypass velocity ($u_{b}$). Using the LMADT-modeled values of $u_{t}$ from Eq. 10, the unconfined freestream velocity, ${U_{\infty}}^{\prime}$, that corresponds to the same $u_{t}$ and dimensional thrust in unconfined flow as for the confined system is given as
${U_{\infty}}^{\prime}=\frac{U_{\infty}\left((u_{t}/U_{\infty})^{2}+C_{T}/4\right)}{u_{t}/U_{\infty}}\ .$ (19)
The unconfined power coefficient (${C_{P}}^{\prime}$), thrust coefficient (${C_{T}}^{\prime}$), and tip-speed ratio ($\lambda^{\prime}$) are then calculated using Eqs. 16, 17 and 18 with ${U_{\infty}}^{\prime}$ in place of $u_{b}$. As shown in Figure 17, up to ${C_{T}}^{\prime}\approx 1$, the ${C_{P}}^{\prime}-\lambda^{\prime}$ and ${C_{T}}^{\prime}-\lambda^{\prime}$ curves exhibit similar collapse across blockages to bluff-body scaling (maximum $\Delta{C_{P}}^{\prime}\approx 0.018$; maximum $\Delta{C_{T}}^{\prime}\approx 0.04$), with the values of ${C_{P}}^{\prime}$ and ${C_{T}}^{\prime}$ generally exceeding those of ${C_{P}}^{b}$ and ${C_{T}}^{b}$ at comparable re-scaled tip-speed. However, for LMADT in unconfined flow, the thrust coefficient cannot exceed unity as this requires reversed flow in the turbine wake. Thus, above the $\lambda^{\prime}$ corresponding to ${C_{T}}^{\prime}\approx 1$, only non-physical solutions exist, and the unconfined efficiency and thrust coefficient curves no longer collapse. Given this constraint on ${C_{T}}^{\prime}$, a key advantage of Maskell-inspired bluff-body scaling over a conventional Glauert-derived scaling is the ability to describe the performance of high-blockage turbines over a wider range of confined $C_{T}$—and thus a wider range of $\beta$ and $\lambda$—since this scaling does not rely on the relationship between $C_{T}$ and $u_{w}$ in unconfined flow.
Since the potential flow model proposed by Steiros et al. [24] also relaxes the constraint on ${C_{T}}^{\prime}$, the question naturally arises as to whether this model—either coupled with Glauert’s theory as in Steiros et al. [24] or with an adaptation of Maskell’s theory following Whelan et al. [2]—can provide a better prediction of flow physics and performance self-similarity for these experiments. As shown in Figure 17, Steiros et al.’s model improves on the weaknesses of traditional blockage corrections at high $\lambda^{\prime}$ while providing similar ${C_{P}}^{\prime}$ and ${C_{T}}^{\prime}$ to that obtained with Houlsby et al.’s model at lower $\lambda^{\prime}$. Furthermore, a Maskell-inspired scaling using the $u_{b}$ calculated from Steiros et al.’s model (Figure 17) achieves a comparable collapse to that based on Houlsby et al.’s model for ${C_{T}}^{b}$ (albeit collapsing to higher values), but somewhat worse collapse for ${C_{P}}^{b}$, particularly at the highest blockages tested. Overall, the scalings in Figure 17 demonstrate the descriptiveness of Maskell-inspired bluff-body scaling for array performance is not exclusive to a particular analytical model type. Notably, despite neglecting free surface deformation, Steiros et al.’s model provides slightly more accurate estimates of the experimental bypass velocities, although the measured wake velocity remains poorly predicted by this model (see supplemental information [59]). However, because Steiros et al.’s model does not provide a materially better scaling for performance, we continue with the open-channel LMADT model of Houlsby et al. for the remainder of this study.
In the present work, measurements of array performance at $\beta\approx 0\%$ are not available, such that the accuracy of Eqs. 16, 17 and 18 for predicting unconfined performance cannot be assessed directly. However, we expect this Maskell-inspired scaling relationship to describe self-similar array performance only under conditions where turbine dynamics resemble those of a bluff body. While the results in this study demonstrate this is likely the case for $\beta=30-55\%$, Araya et al. [70] showed that cross-flow turbine wakes at even lower blockage ($\beta=20\%$) resemble those of a solid cylinder as the “dynamic solidity” of the turbine increases via increasing tip-speed ratio or rotor solidity. Consequently, we hypothesize that the range of blockage ratios over which bluff-body scaling can capture self-similar aspects of array performance depends on the dynamic solidity of the turbine or array. This may explain why the ${C_{T}}^{b}-\lambda^{b}$ curves in Figures 17 and 17 increasingly collapse as $\lambda$ (and thus, $C_{T}$) increases, and is consistent with Whelan et al.’s [2] original framing of Eqs. 17 and 18 as a blockage correction for highly-loaded turbines. Similarly, the poorer agreement between bluff-body scaled performance at $\beta=36\%$ and low-blockage performance at $\beta=3\%$ observed by Ross and Polagye [23] (who used a cross-flow turbine with similar solidity to this study) suggests limitations of bluff body scaling for describing low-blockage or unconfined performance. In these regimes, the blockage correction proposed by Steiros et al. [24] (Figure 17) has clear advantages for high-solidity turbines.
V Prediction of Confined Performance
The self-similar array performance yielded by Maskell-inspired bluff-body scaling provides a pathway for predicting performance at a given blockage ratio using measurements or simulations at another. For example, such information would be useful for describing time-varying performance in a tidal channel with changing water depth and, therefore, blockage. However, only a handful of prior studies have considered “correcting” performance to non-zero blockages. Whelan et al. [2] applied their Maskell-inspired bluff-body blockage correction in reverse to predict the thrust coefficients of a turbine at $\beta=64\%$ from unconfined thrust calculated from BEM, but did not detail the procedure or quantify prediction accuracy. Kinsey and Dumas [17] proposed an empirical forecasting method based on their observation that both the power and thrust coefficients exhibit approximately linear relationships with $\beta$ at constant $\lambda$ (similar to that in Figure 7). Specifically, at several $\lambda$, Kinsey and Dumas develop a linear regression for confined performance from simulations at $\beta=20\%$, and estimated the corresponding performance at $\beta=0\%$ using the LMADT blockage correction of Barnsley and Wellicome [28]. Performance at other $\beta$ was then obtained by interpolation or extrapolation. Although this method was able to accurately predict the power and thrust coefficients at $\beta=10\%$, predictions were less accurate for extrapolation to $\beta\geq 50\%$ where the $C_{T}-\beta$ and $C_{P}-\beta$ relationships are less linear. Finally, the blockage correction developed by Steiros et al. [24] can be used to “correct” turbine performance to arbitrary blockage by simply specifying a non-zero target value of $\beta$ in their equation set. While Steiros et al. explicitly highlight this capability, they do not demonstrate it or evaluate prediction accuracy.
Here, a method for forecasting confined performance based on Houlsby et al.’s LMADT model and the Maskell-inspired bluff-body scaling is described. We expect that this approach is likely similar to that employed by Whelan et al. [2]. A separate evaluation of the forecasting method of Steiros et al. [24] as applied to the experimental data is provided as supplementary material [59]. Consider a turbine in a channel (or an array of equally-spaced turbines spanning the channel width) with known performance at blockage ratio $\beta_{1}$ and Froude number $Fr_{h}$. At a given operating condition (e.g., $\lambda_{1}$) the turbine’s thrust coefficient is $C_{T,1}$ and bypass velocity is $u_{b,1}$. As for Glauert-derived blockage corrections, for the same turbine or array operating at a different blockage ratio, $\beta_{2}$, assume that there is a freestream velocity, $U_{\infty,2}$, that yields the same dimensional thrust on the turbine as at $\beta_{1}$ and $U_{\infty,1}$:
$T_{1}=T_{2}\ .$ (20)
Therefore,
$C_{T,2}=C_{T,1}{\left(\frac{U_{\infty,1}}{U_{\infty,2}}\right)}^{2}\ .$ (21)
Since the thrust on the turbine scales with the bypass velocity in a Maskell-inspired approach, equal thrust at $\beta_{1}$ and $\beta_{2}$ requires equal bypass velocity:
$u_{b,1}=u_{b,2}=u_{b}\ .$ (22)
Therefore, $C_{T,2}$ can be calculated from Eq. 21 if the value of $U_{\infty,2}$ that yields $u_{b}$ at $\beta_{2}$ can be determined. In other words, $C_{T,1}$ and $C_{T,2}$ are connected through a common value of ${C_{T}}^{b}$ when Maskell-inspired bluff-body scaling is applied as
${C_{T}}^{b}=C_{T,1}{\left(\frac{U_{\infty,1}}{u_{b}}\right)}^{2}=C_{T,2}{\left(\frac{U_{\infty,2}}{u_{b}}\right)}^{2}\ .$ (23)
Since rotor power also scales with the bypass velocity, $C_{P,2}$ is given as
$C_{P,2}=C_{P,1}{\left(\frac{U_{\infty,1}}{U_{\infty,2}}\right)}^{3}\ .$ (24)
Finally, assuming constant rotation rate (i.e., $\omega_{1}=\omega_{2}$),
$\lambda_{2}=\lambda_{1}{\left(\frac{U_{\infty,1}}{U_{\infty,2}}\right)}\ .$ (25)
In Eqs. 21 and 24, it is implicitly assumed that $\rho_{1}$ = $\rho_{2}$ (and thus, $\nu_{1}$ = $\nu_{2}$), which implies a difference in the freestream $Re_{D}$ (Eq. 4) between $\beta_{1}$ and $\beta_{2}$. However, the Reynolds number based on the bypass velocity ($u_{b}D/\nu$) remains constant. We note that the inability of LMADT to describe $u_{w}$ and, presumably, $u_{t}$ for relatively high confinement is not a limitation here, since LMADT is able to accurately model the relationship between $\beta$, $C_{T}$, and $u_{b}/U_{\infty}$.
Figure 18: Flowchart describing the procedure for forecasting performance at $\beta_{2}$ based on performance at $\beta_{1}$, assuming constant Froude number between $\beta_{1}$ and $\beta_{2}$. Gray boxes indicate known or specified inputs, yellow circles/ovals indicate equation solving, blue boxes indicate quantities associated with bluff-body scaling, and green boxes indicate the outputs at $\beta_{2}$. The dashed arrows indicate steps where iteration is required.
Using $u_{b}$ as a bridge between $\beta_{1}$ and $\beta_{2}$, LMADT can be used to calculate $U_{\infty,2}$ from known quantities at $\beta_{1}$. The procedure for calculating $U_{\infty,2}$ using the LMADT model of Houlsby et al. [1] is outlined in Figure 18. Here, we assume that the Froude numbers at $\beta_{1}$ and $\beta_{2}$ are constant and subcritical as for the experimental data, although constant Froude number is not required. First, $u_{b}$ is determined at $\beta_{1}$ by numerically solving Eqs. 8 and 9, using $C_{T,1}$, $U_{\infty,1}$, and $Fr_{h}$ as inputs. Next, a reasonable value of $U_{\infty,2}$ is guessed (e.g., $U_{\infty,2}=U_{\infty,1}$), from which a corresponding $C_{T,2}$ is calculated from Eq. 21. Equations 8 and 9 are then solved separately for $u_{w}$ at $\beta_{2}$ using $u_{b}$, $Fr_{h}$, and the guesses for $U_{\infty,2}$ and $C_{T,2}$. The resulting values of $u_{w}$ are compared, and this process is iterated with a new guess for $U_{\infty,2}$ until the value of $U_{\infty,2}$ minimizes the difference between Eq. 8 and Eq. 9. With $U_{\infty,2}$ known, $C_{T,2}$, $C_{P,2}$ and $\lambda_{2}$ may be calculated from $C_{T,1}$, $C_{P,1}$, and $\lambda_{1}$ via Eqs. 21, 24 and 25. Note, however, that $C_{T,1}$ is the only performance metric at $\beta_{1}$ that influences the value of $U_{\infty,2}$.
In the limiting case of $\beta_{2}\rightarrow 0$, $U_{\infty,2}\rightarrow u_{b}$ since the freestream velocity and bypass velocity are equal in unconfined flow. Consequently, Eqs. 21, 24 and 25 revert to Eqs. 16, 17 and 18 corresponding to a Maskell-inspired blockage correction (Figure 17) with the implicit assumption that a bluff-body model is still representative of turbine dynamics at low blockage. Conversely, as $\beta_{2}\rightarrow 1$, the solution space is constrained by the requirements of equal bypass velocity and $Fr_{h}<1$ at $\beta_{1}$ and $\beta_{2}$, such that physically meaningful subcritical solutions at $\beta_{2}$ do not exist for all $\beta_{1}$ and $C_{T,1}$. The maximum value of $\beta_{2}$ at which a physical solution can be obtained depends on $C_{T,1}$, $\beta_{1}$, and the Froude numbers at $\beta_{1}$ and $\beta_{2}$; a visual representation of this limit for the experimental data is provided as supplementary material [59]. However, for the performance data in this study, physically meaningful solutions exist for all practically-achievable blockages (i.e., $\beta_{2}\leq 65\%$).
[…]
Figure 19: 19-19 Predicted thrust coefficient and 19-19 error between the predicted thrust coefficient and measured thrust coefficient ($C_{T,0}$, shown as the gray lines in 19-19) for $\beta_{2}=30.0\%$, $45.0\%$, and $55.0\%$. Color indicates the difference between the blockage ratio of the performance data that was used to make each prediction ($\beta_{1}$) and the target blockage ratio ($\beta_{2}$). Note the different axes limits in 19-19.
[…]
Figure 20: 20-20 Predicted efficiency and 20-20 error between the predicted efficiency and measured efficiency ($C_{P,0}$, shown as the black lines in 20-20) for $\beta_{2}=30.0\%$, $45.0\%$, and $55.0\%$. Color indicates the difference between the nominal blockage ratio of the performance data that was used to make each prediction ($\beta_{1}$) and the target blockage ratio ($\beta_{2}$). Note the different axes limits for each $\beta_{2}$ in 20-20.
To quantify the effectiveness of this forecasting method, the measured array performance at $\beta_{1}=30.0\%-55.0\%$ (Figure 6) is used to predict array performance at $\beta_{2}$, with the resulting $C_{T,2}$ and $C_{P,2}$ at each operating point compared to the experimentally measured performance (denoted as $C_{T,0}$ and $C_{P,0}$). The predicted thrust coefficients at $\beta_{2}=30.0\%$, $45.0\%$ and $55.0\%$ and the corresponding difference between the predicted and measured thrust coefficients are given in Figure 19. The predictions are within $\sim\!0.2$ of measurements at $\beta_{2}=30.0\%$ and $45.0\%$ (Figures 19 and 19) and within $\sim\!0.3$ at $\beta_{2}=55.0\%$ (Figure 19). In general, the magnitude of the thrust prediction error increases with the difference between $\beta_{1}$ and $\beta_{2}$ (i.e., predictions further from the starting blockage are less accurate), and the absolute prediction error tends to grow as $\beta_{2}$ increases for the same difference between $\beta_{1}$ and $\beta_{2}$. In relative terms, these errors are appreciable at lower $\lambda_{2}$ where $C_{T}$ is relatively low, but decline for higher $\lambda_{2}$. Since the foundation of this forecasting method is Maskell-inspired bluff-body scaling (Eqs. 16, 17 and 18), the differences in the predicted $C_{T,2}$ across $\beta_{1}$ are driven by differences in ${C_{T}}^{b}$ across $\beta_{1}$. Thus, the shapes and relative positions of the ${C_{T}}^{b}-\lambda^{b}$ curves across $\beta_{1}$ (Figure 17) dictate the shapes and relative positions of the resulting $C_{T,2}-\lambda_{2}$ curves when scaled to $\beta_{2}$ (Figures 19, 19 and 19). For the current data set, at low to moderate tip-speed ratios this corresponds to underpredicted $C_{T,2}$ if $\beta_{1}>\beta_{2}$ and overpredicted $C_{P,2}$ if $\beta_{1}<\beta_{2}$, except at relatively high $\lambda_{2}$ where this trend inverts.
The corresponding predicted power coefficients and prediction error are shown in Figure 20. The predicted power coefficients are within $\sim\!0.05$ of measurements at $\beta_{2}=30.0\%$ (Figure 20), $\sim\!0.1$ at $45.0\%$ (Figure 20), and $\sim\!0.2$ at $\beta_{2}=55.0\%$ (Figure 20). The greatest prediction error tends to occur near the performance peak and at tip-speed ratios well beyond the optimal operating point. However, the tip-speed ratio corresponding to maximum performance is well-predicted (Figures 20, 20 and 20). Since only the thrust influences the outcome of the bluff-body forecasting procedure, trends in the $C_{P,2}-\lambda_{2}$ curves with $\beta_{1}$ resemble those for the ${C_{T}}^{b}-\lambda^{b}$ curves.
The proposed forecasting method combines linear momentum theory and the Maskell-inspired bluff-body scaling to enable predictions for confined turbine performance from experimental or simulation data. However, since this method utilizes $u_{b}$ as a link between $\beta_{1}$ and $\beta_{2}$, the prediction accuracy depends on whether the bypass velocity is descriptive of array performance at both $\beta_{1}$ and $\beta_{2}$. Consequently, the effectiveness of this forecasting scheme outside of the tested range of $\beta$ is unknown, and the method is likely less accurate when turbine dynamics depart from a bluff-body model. If there exists a transitional condition at which both a Maskell-inspired LMADT scaling (Figure 17) and traditional Glauert-derived LMADT scaling (Figure 17) provide similar results, then a hybrid method that translates between these flow regimes might allow for forecasting performance across a wider ranges of blockages. However, the existence of such a transitional condition is speculative.
VI Conclusions
This research investigates the performance and near-wake flow fields of a cross-flow turbine array at the upper end of practically achievable blockage ratios and demonstrates the efficacy of theoretical frameworks for describing the dynamics of real turbines under these conditions. The performance and near-wake flow field of a two-turbine array were experimentally characterized over a wide range of operating conditions at array blockage ratios from $30\%$ to $55\%$, holding other non-dimensional flow parameters constant to isolate blockage effects. The observed array performance is consistent with trends observed in prior work while providing a uniquely high-resolution view of the manner in which performance evolves with the blockage ratio at relatively high confinements. This highlights relatively subtle trends, such as linear behavior of both the power and thrust coefficients with blockage for low-to-moderate $\beta$ and $\lambda$. Similarly, the measured near-wake velocity fields quantitatively illustrate how the wake and bypass flows change with $\beta$ and $\lambda$. This work also provides an openly-available set of experimental performance and flow-field data for cross-flow turbines at high blockage, which will aid in the validation of numerical simulations and reduced-order models that seek to examine confinement effects and inform the design of full scale systems that can exploit these physics.
A key outcome of this work is the demonstration that linear momentum actuator disk models can quantitatively describe aspects of real turbine arrays in highly confined flows. It is shown that, at a location $0.6$ turbine diameters downstream of the array, the velocity of the flow bypassing the array measured using PIV is well represented by the bypass velocity predicted from the open-channel LMADT model of Houlsby et al. [1] using the measured array thrust coefficients and inflow conditions as inputs. This result was confirmed for a wide range of blockages and tip-speed ratios, with the bypass velocity calculated from LMADT within approximately $8\%$ of the measured velocity between the turbines at the array centerline at all conditions. The residual deviation between modeled and measured bypass velocities may be a consequence of near-wake mixing that is neglected in LMADT. Following Whelan et al. [2] and inspired by the bluff-body theory of Maskell [3], scaling the array power and thrust coefficients by this analytical bypass velocity collapses these coefficients across the tested blockage and tip-speed ratios. These results highlight similar dynamics between high-blockage turbine arrays and bluff bodies, and demonstrate that, despite their inherent simplicity, LMADT models can effectively model aspects of the flow around non-ideal turbines in confined flows and provide insights into the salient hydrodynamics that govern turbine performance. Leveraging the connection between array performance and the bypass velocity, an analytical forecasting method based on LMADT is presented by which known turbine performance at one confinement ($\beta_{1}$) can be used to predict turbine performance at a different confinement ($\beta_{2}$) with moderate accuracy. This forecasting method provides a pathway for estimating turbine performance under confined conditions based on the dominant dynamics of high-blockage arrays.
Although this work focuses on cross-flow turbines, the characteristic fluid dynamics of confinement are not exclusive to this archetype, and the analytical models employed are agnostic to turbine geometry. However, the design of the rotor is expected to influence how these characteristic dynamics change with blockage and the tip-speed ratio [13], and the effectiveness of Maskell-inspired bluff-body scaling and forecasting are likely dependent on turbine geometry and operating condition. It is recommended that future work explore the broader application of these analytical methods to turbines at lower confinement, as well as to different turbine designs. Additionally, although this work focuses on the time-averaged array performance and near-wake flow fields, the unsteady fluid dynamics of cross-flow turbines also inform their performance and the properties of the near-wake. Future research should examine how the phase-resolved performance of cross-flow turbines develops with confinement, and how these dynamics influence the temporal and spatial evolution of the flow field in the near- and far-wake regions.
Acknowledgements
This work was supported by the United States Advanced Research Projects Agency – Energy (ARPA-E) under award number DE-AR0001441. The authors would like to thank Gregory Talpey, Gemma Calandra, and Corey Crisp for their assistance in commissioning the high-confinement test-rig, as well as help with data collection. Abigale Snortland, Hannah Ross, Jennifer Franck, and Mukul Dave are acknowledged for insightful conversations regarding the results in this work. We are grateful to the Alice C. Tyler Charitable Trust for upgrades to the experimental flume, including closed loop heating and cooling.
Data Availability
The data that support the findings of this study are openly available in the Dryad data repository at https://doi.org/10.5061/dryad.dfn2z35b5 [72] (DOI pending activation upon acceptance). The MATLAB code used to implement the models of Houlsby et al. [1] and Steiros et al. [24] is available on GitHub at https://github.com/aidan-hunt/turbine-confinement-models.
Supplementary Material
See supplementary material at [URL will be inserted by publisher] for lateral force measurements, estimates of hydraulic efficiency, application of Steiros et al.’s model to the experimental data, a visualization of the solution space for the bluff-body forecasting method, and additional figures of PIV flow fields at $0.6D$ and $1.5D$ downstream of the array.
Author contributions (CRediT)
A. Hunt: Conceptualization, Investigation, Validation, Methodology, Formal Analysis, Visualization, Software, Writing - Original Draft and Review & Editing. A. Athair: Investigation, Validation, Methodology, Formal Analysis, Visualization, Software, Writing - Original Draft and Review & Editing. O. Williams: Conceptualization, Methodology, Supervision, Writing - Review & Editing B. Polagye: Conceptualization, Methodology, Supervision, Writing - Original Draft and Review & Editing, Resources, Funding Acquisition.
Appendix A Influence of Submergence Depth on Performance
[…]
Figure 21: Change in the 21-21 $C_{P}-\lambda$ and 21-21 $C_{T}-\lambda$ curves as a function of $s/h$ and $\lambda$ relative to the baseline performance curves in Figure 6 for $\beta=36.7\%$, $45.0\%$, and $55.0\%$, with general regions of significant ventilation annotated. The value of $s/h$ at each $\beta$ that corresponds to baseline experimental results is indicated in the colorbar and corresponds to the zero line in each tile.
As mentioned in Section II.2, $s/h$ was maximized at each blockage to minimize the impact of ventilation on array performance, particularly at the highest blockage ratios. However, this resulted in variation of $s/h$ across the tested $\beta$ (Table 1). To assess the sensitivity of array performance to submergence depth, the array was also tested at a range of $s/h$ at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ through variation of the rotors’ vertical position in the water column. The ADV position was likewise adjusted such that $U_{\infty}$ was always sampled at the turbine midplane.
The difference between the $C_{P}\!-\!\lambda$ and $C_{T}\!-\!\lambda$ curves presented in Figure 6 and those obtained at different $s/h$ are shown in Figure 21. At $\beta=36.7\%$ and $\beta=45.0\%$, both $C_{P}$ and $C_{T}$ increase slightly as $s/h$ decreases (i.e., the turbines are moved closer to the free surface), particularly for $\lambda$ beyond the performance peak. However, at all $\beta$, once $s/h$ drops below a critical value, ventilation begins, resulting in decreased $C_{P}$ and $C_{T}$ relative to deeper submergence depths. As $\beta$ increases, the onset of ventilation occurs at larger $s/h$ (i.e., deeper submergence), and at a given $\beta$, as $s/h$ decreases, ventilation begins at lower $\lambda$. However, other than ventilation, the effects of submergence depth on both $C_{P}$ and $C_{T}$ are minor relative to the influence of $\beta$ and the associated effects are mainly limited to $\lambda$ beyond the performance peak. Consequently, variation in $s/h$ across the tested $\beta$ is unlikely to significantly affect the trends presented in Figure 6 or the flow field structure.
[…]
Figure 22: Change in the time-averaged normalized streamwise water depth profiles as a function of $s/h$ at 10 $\beta=36.7\%$, $\beta=45.0\%$, and 10 $\beta=55.0\%$ for $2$ $\lambda=1.6$, $2$ $\lambda=2.6$, and $2$ $\lambda=3.6$. The farthest upstream measurement is taken at the center of the channel, whereas depth measurements in the vicinity of the array are directly upstream and downstream of Turbine B (Figure 3). For each case, depth measurements at each streamwise location ($h_{x}$) are normalized by the far-upstream water depth ($h$). The dashed lines indicate the swept area of Turbine B. The value of $s/h$ for each $\beta$ that corresponds to baseline experimental results (Figure 10) is indicated in the colorbar and corresponds to the zero line in each tile.
The normalized submergence depth also slightly affects free surface deformation in the vicinity of the array, as shown by the change in the water depth profiles in Figure 22 for $\lambda=1.6$, $2.6$, and $3.6$ at $\beta=36.7\%$, $45.0\%$, and $55.0\%$. Particularly for $\beta=36.7\%$ and $45.0\%$, decreasing $s/h$ somewhat elevates the free surface upstream of Turbine B and lowers the free surface directly downstream at $x/D=0.5$. In other words, the magnitude of the free surface drop across Turbine B somewhat increases with greater proximity of the rotor to the free surface, and this drop becomes more pronounced as $\lambda$ and $\beta$ increase. This is consistent with the non-ventilated power and thrust trends shown in Figure 21. Trends in free surface deformation with $s/h$ at $\beta=55.0\%$ are less clear, likely due to the significant ventilation that occurs this blockage ratio. Similarly, for all cases trends in the free surface with $s/h$ further downstream ($x/D=1-2$) are more ambiguous, and may be obscured by the relatively low spatial and temporal resolution of the free surface transducers utilized in this study.
Supplemental information for “Experimental validation of a linear momentum and bluff-body model for high-blockage cross-flow turbine arrays”
Aidan Hunt1$\ast$, Ari Athair2, Owen Williams2, and Brian Polagye1
1 Department of Mechanical Engineering, University of Washington.
2 Department of Aeronautics and Astronautics, University of Washington.
$\ast$: To whom correspondence should be addressed: ahunt94@uw.edu.
This PDF file contains the following supplementary information:
  • Lateral force coefficients for the cross-flow turbine array,
  • Estimations of hydraulic efficiency for the cross-flow turbine array,
  • Evaluation of the analytical model of Steiros et al. [24],
  • Solution space for the bluff-body forecasting method based on the model of Houlsby et al. [1],
  • Additional figures of PIV flow fields $0.6D$ and $1.5D$ downstream of the array.
Lateral force coefficients
Since the turbines in the array are counter-rotated with zero phase offset between them, the net lateral force on the array is approximately zero since the array is symmetric about its centerline. However, by defining the directions of positive $L_{A}$ and $L_{B}$ with this counter-rotation in mind as in Figure 3, the array-average lateral force coefficient is nonzero and represents the average lateral force coefficient experienced by an individual rotor in the array as
$C_{L,\mathrm{rotor}}=\frac{L_{A}+L_{B}}{\frac{1}{2}\rho U_{\infty}^{2}A_{\mathrm{turbines}}}\ \ .$ (S.1)
Figure S.1 shows the effect of blockage on $C_{L,\mathrm{rotor}}-\lambda$ curves across the tested $\beta$ and $\lambda$. As for $C_{P}$ and $C_{T}$, the magnitude of $C_{L,\mathrm{rotor}}$ at a particular $\lambda$ tends to increase with the blockage ratio. Additionally in a similar manner to the array power and thrust coefficients, $C_{L,\mathrm{rotor}}$ also exhibits invariance to $\beta$ at lower tip-speed ratios ($\lambda\leq 1$).
Figure S.1: Time-averaged rotor-average lateral force coefficient as a function of $\beta$ and $\lambda$. The shaded regions indicate the interquartile range of the array- and cycle-averaged performance at each $\beta$ and $\lambda$ (the vertical span of the shaded region at each point is similar to the size of the markers for most operating conditions).
Estimating Hydraulic Efficiency
As noted in Section III, the conventional definition of the array power coefficient (Eq. 6) neglects the potential energy that may be harnessed by the array in a confined flow. Consequently, at higher blockage, array efficiency may be better described using a hydraulic efficiency ($\eta$) that considers both the static and dynamic head in the flow:
$\eta=\frac{Q_{A}\omega_{A}+Q_{B}\omega_{B}}{\dot{m}g\Delta h}\ \ ,$ (S.2)
where $\dot{m}$ is the mass flow rate through the channel and $\Delta h$ is the difference in the water depth between far upstream and far downstream. Similar definitions of efficiency were employed by Takamatsu et al. [34] and McAdam et al. [36] in their studies of high-blockage cross-flow turbines. However, both $\dot{m}$ and $\Delta h$ can be difficult to measure in practice. In this study, instrumentation to measure the volumetric flow rate through the flume was unavailable, and the volumetric flow rate is expected to vary with $\lambda$ at each $\beta$ due to interactions between the array and the channel [44]. Additionally, accurate measurements of $\Delta h$ require a channel of sufficient length to allow the flow to fully mix downstream of the array and sufficient measurement resolution to identify this condition. For the experimental set-up in this study, the finite length of the flume downstream of the cross-flow turbine array is likely insufficient to allow the flow to fully mix, and the furthest downstream measurement of the water depth was taken only $2D$ downstream of the array.
Although precise measurements of $\dot{m}$ and $\Delta h$ are not available in this study, we present a rough estimate of $\eta$ for the experimental array across the tested range of $\beta$ and $\lambda$ using the available velocity and depth measurements. Assuming a uniform velocity profile in the Tyler flume, we estimate $\dot{m}\approx\rho U_{\infty}hw$. $\Delta h$ is estimated as the difference between the furthest upstream measured water depth (at $x/D=-5.8$) and furthest downstream measured water depth (at $x/D=2$). The resulting hydraulic efficiencies are expressed as a function of $\beta$ and $\lambda$ in Figure S.2. Unlike $C_{P}$, the values of $\eta$ are $\leq 1$, with the exception of several spikes in the estimated $\eta$ at lower $\lambda$. These spikes occur when $\Delta h\approx 0$ between the two sample locations resulting from an irregular and unsteady free surface, and emphasizes the spatial and temporal resolution required to characterize the free surface deformation for the purposes of computing hydraulic efficiency. As $\lambda$ increases, the thrust on the array results in a more defined free surface profile and less noise in the estimates of hydraulic efficiency, with trends in $\eta$ with blockage largely following those of $C_{P}$ (Figure S.2). Nonetheless, the $\eta$ shown are likely lower than the true hydraulic efficiencies, due to overestimation of the mass flow rate and total head drop from the available velocity and depth measurements.
[…]
Figure S.2: S.2 Time-average hydraulic efficiency estimated for the array, compared to S.2 the time-average array-average $C_{P}$ (reproduced from Figure 6). The axes limits in Figure S.2 are cropped to focus on non-spurious values of $\eta$.
Application of Steiros et al.’s model to the experimental data
S.0.1 Prediction of Near-Wake Velocities
Here, we apply a confinement model developed by Steiros et al. [24] based on two-dimensional potential flow theory to the experimental data to predict the bypass velocities ($u_{b}$) and wake velocities ($u_{w}$) from measurements of $C_{T}$ and $U_{\infty}$ at each $\beta$. In this model, free surface deformation is assumed to be negligible (i.e., $Fr_{h}\rightarrow 0$). Key details of the modeling approach and resulting equation set are summarized here; a detailed derivation is provided in Steiros et al. [24], with the underlying porous plate theory described by Steiros and Hultmark [30]. Steiros et al. follow the theory of Taylor [32] for flow through and around a porous plate by representing the turbine as a line of evenly distributed potential sources in a freestream flow. The induced flow field through and around the turbine is then determined through superposition, with adjustments applied to ensure continuity through the turbine [33] and to allow the wake pressure to fall below the ambient pressure [24]. Flow confinement is imposed through the method of mirror images by considering an infinite number of turbines arranged side-by-side with a distance $w$ between their centers. The thrust on the turbine is then computed by applying conservation of mass, momentum, and energy in a control volume encompassing the modeled flow field around the turbine, yielding
$C_{T}=\frac{4(u^{\ast}-1)(1-u^{\ast})}{(1-\beta)(2-u^{\ast}-u^{\ast}\beta)}\left(\frac{1-u{\ast}}{3}-\frac{1-2u^{\ast}\beta+\beta}{1-\beta}\right),$ (S.3)
where $u^{\ast}=u_{t}/U_{\infty}$ is the normalized velocity through the turbine rotor.
As for the LMADT model of Houlsby et al. [1], in applying the model of Steiros et al. [24] we treat the array as a single turbine with the same blockage ratio as the array. Using the measured $C_{T}$ and $\beta$, Eq. S.3 may be solved numerically for $u^{\ast}$. Then, the bypass velocity and wake velocity may be calculated respectively as
$u_{b}=U_{\infty}\left(\frac{1-2u^{\ast}\beta+\beta}{1-\beta}\right)\ ,$ (S.4)
and
$u_{w}=U_{\infty}\left(\frac{u^{\ast}(1-2u^{\ast}\beta+\beta)}{2-u^{\ast}-u^{\ast}\beta}\right)\ .$ (S.5)
The resulting $u_{b}$ and $u_{w}$ are normalized by the freestream velocity and plotted in Figure S.3. Trends in the near-wake velocities predicted by the model of Steiros et al. [24] are the same as those for the near-wake velocities predicted by Houlsby et al. [1]: $u_{b}/U_{\infty}$ increases with $\lambda$ whereas $u_{w}/U_{\infty}$ decreases with $\lambda$, and both $u_{b}/U_{\infty}$ and $u_{w}/U_{\infty}$ increase with $\beta$. While both models yield similar predictions for the wake and bypass velocities at low $\lambda$, Steiros et al.’s model predicts lower $u_{b}$ and $u_{w}$ than Houlsby et al.’s model as $\lambda$ increases. Despite these differences, the $u_{b}$ obtained from Steiros et al.’s model shows good agreement with the streamwise velocity measured in the bypass region $0.6D$ downstream of the array (Figures S.4, S.4 and S.4, Figure S.5), and is more accurate than that obtained via Houlsby et al.’s model. In contrast, the experimentally measured wake velocity is still poorly predicted by Steiros et al.’s model at this location Figures S.4 and S.4. Flow fields at other $\beta$, $\lambda$, and streamwise locations normalized by $u_{b}$ and $u_{w}$ obtained from Steiros et al.’s model are subsequently provided in Figures S.15, S.16, S.20 and S.21.
[…]
Figure S.3: The S.3 bypass velocities, and S.3 wake velocities (all normalized by the freestream velocity) predicted by the model of Steiros et al. [24] from measurements of array thrust and flow conditions. The near-wake velocities predicted by the LMADT model of Houlsby et al. [1] from Figure 12 are shown in gray for comparison.
[…]
Figure S.4: Time-averaged streamwise velocities measured $0.6D$ downstream of the array at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ and optimal $\lambda$, normalized by the corresponding 13-13 bypass velocity predicted by the model of Steiros et al. [24], and 13-13 wake velocity predicted by the model of Steiros et al. [24]. The black rectangles indicate the projected area of the turbines in the field of view.
[…]
Figure S.5: Time-averaged streamwise velocity profiles measured $0.6D$ downstream of the array normalized by the bypass velocity predicted from the model of Steiros et al. [24] at 14 $\beta=36.7\%$, 14 $45.0\%$, and 14 $55.0\%$ and various $\lambda$. The profiles shown correspond to the average velocity along the middle $70\%$ of the turbine blade span. Vertical gray rectangles correspond to the bypass region.
S.0.2 Performance Scaling
As in Section IV.4, two methods for scaling the array performance based on the model of Steiros et al. [24] are considered. First, a Maskell-inspired bluff-body blockage correction is performed by applying Eqs. 16, 17 and 18 to the measured $C_{P}$, $C_{T}$, and $\lambda$ using the $u_{b}$ computed from Eq. S.4. The resulting ${C_{T}}^{b}-\lambda^{b}$ and ${C_{P}}^{b}-\lambda^{b}$ curves are shown in Figure S.6 and are compared to those obtained using the $u_{b}$ calculated from the LMADT model of Houlsby et al. [1]. Both models yield qualitatively similar scalings, with Steiros et al.’s model generally resulting in higher values of ${C_{T}}^{b}$ and ${C_{P}}^{b}$ as $\lambda^{b}$ increases. While both models highlight self-similarity in $C_{P}$ across blockage ratios, the ${C_{P}}^{b}-\lambda^{b}$ curves collapse more tightly for Houlsby et al.’s model.
[…]
Figure S.6: Comparison of the S.6 ${C_{T}}^{b}$ and S.6 ${C_{P}}^{b}$ obtained from the measured array performance (Figure 6) when different analytical models are used to estimate $u_{b}$. The curves shown are the same as in Figures 17 and 17.
Second, a Glauert-derived blockage correction is performed using the model of Steiros et al. [24] by determining the unconfined freestream velocity, ${U_{\infty}}^{\prime}$, that results in the same $u_{t}$ and dimensional thrust at $\beta=0\%$ as for the confined condition, i.e.
$u_{t}={u_{t}}^{\prime},$ (S.6)
$T={T}^{\prime}.$ (S.7)
Therefore, it follows that
${C_{T}}^{\prime}({U_{\infty}}^{\prime})^{2}={C_{T}}(U_{\infty})^{2}\ ,$ (S.8)
and consequently,
${C_{T}}^{\prime}={C_{T}}\left(\frac{U_{\infty}}{{U_{\infty}}^{\prime}}\right)^{2}={C_{T}}\left(\frac{u^{\ast}}{{u^{\ast}}^{\prime}}\right)^{2}$ (S.9)
since $u_{t}$ = ${u_{t}}^{\prime}$. Following the procedure outlined by Steiros et al. [24], the blockage correction is performed by first calculating $u^{\ast}$ from the known $C_{T}$ and $\beta$ at a given confined condition using Eq. S.3. Then, ${u^{\ast}}^{\prime}$ is calculated by simultaneously solving Eqs. S.3 and S.9 for the value of ${u^{\ast}}^{\prime}$ that yields equivalent values of ${C_{T}}^{\prime}$, with $\beta=0\%$ used in Eq. S.3. Once ${u^{\ast}}^{\prime}$ is known, ${C_{T}}^{\prime}$ is computed from Eq. S.9 and ${C_{P}}^{\prime}$ and $\lambda^{\prime}$ are computed as
${C_{P}}^{\prime}={C_{P}}\left(\frac{u^{\ast}}{{u^{\ast}}^{\prime}}\right)^{3}\,$ (S.10)
${\lambda}^{\prime}={\lambda}\left(\frac{u^{\ast}}{{u^{\ast}}^{\prime}}\right)$ (S.11)
[…]
Figure S.7: Comparison of the S.7 ${C_{T}}^{\prime}$ and S.7 ${C_{P}}^{\prime}$, obtained from the measured array performance (Figure 6) when different analytical models are used to estimate ${U_{\infty}}^{\prime}$. The curves shown are the same as in Figures 17 and 17.
The resulting ${C_{T}}^{\prime}-\lambda^{\prime}$ and ${C_{P}}^{\prime}-\lambda^{\prime}$ curves are shown in Figure S.7 and are compared to those obtained using $U_{\infty}^{\prime}$ calculated from the LMADT model of Houlsby et al. [1]. Blockage corrections based on both models generally agree up to $\lambda^{\prime}\approx 1.25$, where ${C_{T}}^{\prime}\approx 0.9$. However, Steiros et al.’s model produces physically-valid values of ${C_{T}}^{\prime}$ and ${C_{P}^{\prime}}$ beyond the $\lambda^{\prime}$ corresponding to ${C_{T}}^{\prime}=1$, whereas physically-valid solutions to Houlsby et al.’s model are restricted to ${C_{T}}^{\prime}\leq 1$ as expected. While the ${C_{T}}^{\prime}-\lambda^{\prime}$ and ${C_{P}}^{\prime}-\lambda^{\prime}$ curves obtained using the Steiros et al.’s model generally collapse across $\beta$, the curves tend to diverge more at higher $\lambda^{\prime}$. This result is consistent with observations by Steiros et al. in their validation of this model using measurements of drag on flat plates with various porosities and blockage ratios. Specifically, Steiros et al. noted that the model accuracy decreased with increasing blockage ratio and decreasing plate porosity (analogous to increasing tip-speed ratio here).
S.0.3 Performance Forecasting
The Glauert-derived blockage correction method of Steiros et al. [24] may also used to “correct” confined performance performance at one blockage ratio ($\beta_{1}$) to a different blockage ratio ($\beta_{2}$) by simply specifying a non-zero target value of $\beta$ when solving Eq. S.3 for ${u^{\ast}}^{\prime}$. While Steiros et al. note this forecasting capability, they did not demonstrate it or evaluate the accuracy of such forecasts in their study. Here, we apply this forecasting approach to the experimental data and present the results in a similar manner to Section V. Figure S.8 shows the predicted thrust coefficients at $\beta_{2}=30.0\%$, $45.0\%$, and $55.0\%$ when Steiros et al.’s forecasting method is applied to the experimental data. $C_{T}$ is accurately predicted at lower $\lambda$, but error between the predicted and measured $C_{T}$ becomes large at higher $\lambda$. Similar trends are observed for the predicted $C_{P}$ using this method (Figure S.9). For both $C_{T}$ and $C_{P}$, forecast accuracy becomes worse for a larger difference between $\beta_{1}$ and $\beta_{2}$. This result is consistent with Steiros et al.’s observations of the limitations of their model’s accuracy at high $\lambda$. Similar to the LMADT-based bluff-body forecast method evaluated in Section V, trends in prediction accuracy resemble trends in blockage-corrected ${C_{T}}^{\prime}-\lambda^{\prime}$ and ${C_{P}}^{\prime}-\lambda^{\prime}$ curves Figure S.7, which form the basis for confined performance forecasting based in Glauert’s theory.
[…]
Figure S.8: Confined $C_{T}$ forecasted from the experimental data using the Glauert-derived blockage correction method proposed by Steiros et al. [24]. S.8-S.8 Predicted thrust coefficient and S.8-S.8 error between the predicted thrust coefficient and measured thrust coefficient ($C_{T,0}$, shown as the gray lines in S.8-S.8) for $\beta_{2}=30.0\%$, $45.0\%$, and $55.0\%$. Color indicates the difference between the blockage ratio of the performance data that was used to make each prediction ($\beta_{1}$) and the target blockage ratio ($\beta_{2}$). Note the different axes limits in S.8-S.8.
[…]
Figure S.9: Confined $C_{P}$ forecasted from the experimental data using Glauert-derived the blockage correction method proposed by Steiros et al. [24]. S.9-S.9 Predicted power coefficient and S.9-S.9 error between the predicted power coefficient and measured power coefficient ($C_{T,0}$, shown as the gray lines in S.9-S.9) for $\beta_{2}=30.0\%$, $45.0\%$, and $55.0\%$. Color indicates the difference between the blockage ratio of the performance data that was used to make each prediction ($\beta_{1}$) and the target blockage ratio ($\beta_{2}$). Note the different axes limits in S.9-S.9.
While Figures S.8 and S.9 demonstrate a Glauert-derived performance forecasting method based upon the model of Steiros et al. [24], the collapse in the ${C_{T}}^{b}-\lambda^{b}$ curves in Figure S.6 suggests that a bluff-body forecast based on this model may also be suitable. Here, we propose a bluff-body forecast method based on Steiros et al.’s model by following a similar procedure as in Section V. We assume that there is some freestream velocity $U_{\infty,2}$ at $\beta_{2}$ that yields the same dimensional thrust and bypass velocity at $\beta_{2}$ as at $\beta_{1}$:
$u_{b,1}=u_{b,2}=u_{b}\ ,$ (S.12)
$T_{1}=T_{2}=T\ .$ (S.13)
Therefore, as in Equation 21,
$C_{T,2}=C_{T,1}{\left(\frac{U_{\infty,1}}{U_{\infty,2}}\right)}^{2}\ .$ (S.14)
To forecast performance at $\beta_{2}$ from known performance at $\beta_{1}$, a reasonable value of $U_{\infty,2}$ is guessed. Then, using this guess and the known $u_{b}$ (which is constant between $\beta_{1}$ and $\beta_{2}$), Equation S.4 is solved for $u^{\ast}_{2}$ at $\beta_{2}$. $C_{T,2}$ is then solved from Equation S.3 using $\beta_{2}$ and $u^{\ast}_{2}$, and this value is compared to the $C_{T,2}$ obtained from Equation S.14. This process is iterated until the difference between the values of $C_{T,2}$ obtained from Equations S.3 and S.14 is minimized. Finally, with $U_{\infty,2}$ known, $C_{T,2}$ is obtained from Equation S.14, and $C_{P,2}$ and $\lambda_{2}$ are respectively calculated as
$C_{P,2}=C_{P,1}{\left(\frac{U_{\infty,1}}{U_{\infty,2}}\right)}^{3}\ ,$ (S.15)
$\lambda_{2}=\lambda_{1}{\left(\frac{U_{\infty,1}}{U_{\infty,2}}\right)}\ .$ (S.16)
[…]
Figure S.10: Confined $C_{T}$ forecasted from the experimental data using a bluff-body method and the model of Steiros et al. [24]. S.10-S.10 Predicted thrust coefficient and S.10-S.10 error between the predicted thrust coefficient and measured thrust coefficient ($C_{T,0}$, shown as the gray lines in S.10-S.10) for $\beta_{2}=30.0\%$, $45.0\%$, and $55.0\%$. Color indicates the difference between the blockage ratio of the performance data that was used to make each prediction ($\beta_{1}$) and the target blockage ratio ($\beta_{2}$). Note the different axes limits in S.10-S.10.
[…]
Figure S.11: Confined $C_{P}$ forecasted from the experimental data using a bluff-body method and the model of Steiros et al. [24]. S.11-S.11 Predicted thrust coefficient and S.11-S.11 error between the predicted thrust coefficient and measured thrust coefficient ($C_{P,0}$, shown as the gray lines in S.11-S.11) for $\beta_{2}=30.0\%$, $45.0\%$, and $55.0\%$. Color indicates the difference between the blockage ratio of the performance data that was used to make each prediction ($\beta_{1}$) and the target blockage ratio ($\beta_{2}$). Note the different axes limits in S.11-S.11.
Figures S.10 and S.11 show the thrust coefficients and power coefficients, respectively, predicted at $\beta_{2}=30.0\%$, $45.0\%$, and $55.0\%$ when a bluff-body forecast based on Steiros et al.’s model is applied to the experimental data. As for the Maskell-inspired bluff-body forecasts based on the LMADT model of Houlsby et al. [1] presented in Figures 19 and 20, trends in forecasting accuracy follow trends in the ${C_{T}}^{b}-\lambda^{b}$ and ${C_{P}}^{b}-\lambda^{b}$ curves, which serve as the basis for bluff-body forecasts. Bluff-body prediction accuracy for $C_{T}$ using Steiros et al.’s model is comparable to that of Houlsby et al.’s model, owing to the similar collapse in the ${C_{T}}^{b}-\lambda^{b}$ curves obtained by both models (Figure S.6). However, bluff-body prediction accuracy for $C_{P}$ for Steiros et al.’s model is worse than for Houlsby et al.’s model—especially at high $\lambda$—due to the worse collapse in the ${C_{P}}^{b}-\lambda^{b}$ observed (Figure S.6). As for all other forecasts considered in this study, forecast accuracy for both $C_{T}$ and $C_{P}$ becomes worse for a larger difference between $\beta_{1}$ and $\beta_{2}$, as well as at higher $\beta_{2}$ for a given difference between $\beta_{1}$ and $\beta_{2}$.
Comparing the two performance forecasting methods based on Steiros et al.’s model examined here, a Glauert-derived forecast yields accurate predictions of $C_{T}$ at lower $\lambda$, whereas a bluff-body forecast for $C_{T}$ is more accurate at higher $\lambda$. Therefore, a hybrid approach that utilizes either a Glauert-derived forecast or a bluff-body forecast depending on the regime could provide adequate prediction accuracy over all $\lambda$. However, such a hybrid approach would not be effective for predicting $C_{P}$, as both Glauert-derived forecasts and bluff-body forecasts of $C_{P}$ using Steiros et al.’s model become less accurate as $\lambda$ increases.
Solution Space for LMADT-based Bluff-body Forecasting
Figure S.12: Conditions under which physically meaningful predictions for performance at $\beta_{2}$ can be obtained using the bluff-body forecasting method based on Houlsby et al.’s [1] model described in Section V and the collected performance data (Figure 6). Here, it is assumed that $Fr_{h}$ is constant between $\beta_{1}$ and $\beta_{2}$. The white regions in each panel represent combinations of $C_{T,1}$ and $\beta_{2}$ that yield physical, sub-critical solutions, whereas the black regions indicate combinations that yield only non-physical solutions. The gray regions in each panel indicate $C_{T,1}$ outside of the range of values measured at each $\beta_{1}$, and the dashed blue line in each panel indicates $\beta_{1}$ relative to the range of $\beta_{2}$ that can be forecasted.
Additional PIV Flow Field Figures
Figure S.13: Time-averaged streamwise velocity $0.6D$ downstream of the array as function of $\beta$ and $\lambda$, normalized by the corresponding bypass velocity, $u_{b}$, predicted by the model of Houlsby et al. [1]. The black rectangles in each tile indicate the projected area of the turbines in the field of view. Tiles corresponding to the optimal $\lambda$ for each blockage are outlined in purple (tiles (d), (h), and (k)), and correspond to Figures 13, 13 and 13.
Figure S.14: Time-averaged streamwise velocity $0.6D$ downstream of the array as function of $\beta$ and $\lambda$, normalized by the corresponding wake velocity, $u_{w}$, predicted by the model of Houlsby et al. [1]. The black rectangles in each tile indicate the projected area of the turbines in the field of view. Flow fields corresponding to the optimal $\lambda$ for each blockage are outlined in purple (tiles (d), (h), and (k)), and correspond to Figures 13, 13 and 13.
Figure S.15: Time-averaged streamwise velocity $0.6D$ downstream of the array as function of $\beta$ and $\lambda$, normalized by the corresponding bypass velocity, $u_{b}$, predicted by the model of Steiros et al. [24]. The black rectangles in each tile indicate the projected area of the turbines in the field of view. Tiles corresponding to the optimal $\lambda$ for each blockage are outlined in purple (tiles (d), (h), and (k)), and correspond to Figures S.4, S.4 and S.4.
Figure S.16: Time-averaged streamwise velocity $0.6D$ downstream of the array as function of $\beta$ and $\lambda$, normalized by the corresponding wake velocity, $u_{w}$, predicted by the model of Steiros et al. [24]. The black rectangles in each tile indicate the projected area of the turbines in the field of view. Flow fields corresponding to the optimal $\lambda$ for each blockage are outlined in purple (tiles (d), (h), and (k)), and correspond to Figures S.4, S.4 and S.4.
Figure S.17: Time-averaged streamwise velocity $1.5D$ downstream of the array as function of $\beta$ and $\lambda$. Velocity measurements are normalized by the freestream velocity measured by the ADV. The black rectangles in each tile indicate the projected area of the turbines in the field of view. Flow fields corresponding to the optimal $\lambda$ for each blockage are outlined in purple (tiles (d), (h), and (k)).
Figure S.18: Time-averaged streamwise velocity $1.5D$ downstream of the array as function of $\beta$ and $\lambda$, normalized by the corresponding bypass velocity, $u_{b}$, predicted by the model of Houlsby et al. [1]. The black rectangles in each tile indicate the projected area of the turbines in the field of view. Tiles corresponding to the optimal $\lambda$ for each blockage are outlined in purple (tiles (d), (h), and (k)).
Figure S.19: Time-averaged streamwise velocity $1.5D$ downstream of the array as function of $\beta$ and $\lambda$, normalized by the corresponding wake velocity, $u_{w}$, predicted by the model of Houlsby et al. [1]. The black rectangles in each tile indicate the projected area of the turbines in the field of view. Flow fields corresponding to the optimal $\lambda$ for each blockage are outlined in purple (tiles (d), (h), and (k)).
Figure S.20: Time-averaged streamwise velocity $1.5D$ downstream of the array as function of $\beta$ and $\lambda$, normalized by the corresponding bypass velocity, $u_{b}$, predicted by the model of Steiros et al. [24]. The black rectangles in each tile indicate the projected area of the turbines in the field of view. Tiles corresponding to the optimal $\lambda$ for each blockage are outlined in purple (tiles (d), (h), and (k)).
Figure S.21: Time-averaged streamwise velocity $1.5D$ downstream of the array as function of $\beta$ and $\lambda$, normalized by the corresponding wake velocity, $u_{w}$, predicted by the model of Steiros et al. [24]. The black rectangles in each tile indicate the projected area of the turbines in the field of view. Flow fields corresponding to the optimal $\lambda$ for each blockage are outlined in purple (tiles (d), (h), and (k)).